### v1

# Eigenvalue Gaps of the Laplacian of Random Graphs

Abstract.
We show that, with very high probability, the random graph Laplacian has simple spectrum. Our method provides a quantitatively effective estimate of the spectral gaps. Along the way, we establish results on affine no-gaps delocalization, no-structure delocalization, overcrowding and small entries of the eigenvectors for the Laplacian model. These findings are of independent interest.
Part of this work was completed while the third author was supported by a Simons Fellowship.
1. Introduction
Let $G_{n}=G(n,p)$ be a random Erdős-Rényi graph with a fixed parameter $p$, and let $A_{n}$ be its adjacency matrix. Partly motivated by a question by Babai (related to the notorious graph isomorphism problem in theoretical computer science), it has been shown in [TV17, NTV17] that with high probability the spectrum sequence $\lambda_{1}(A_{n})\leq\dots\leq\lambda_{n}(A_{n})$ of $A_{n}$ is simple. More specifically
Theorem 1.1.
[NTV17, Theorem 2.7] Let $10<p<$ 1 be fixed and let $A$ be any constant. Then for any $\delta>n^{-A}$, we have
$\max_{1\leq i\leq n-1}{\mathbb{P}}(|\lambda_{i+1}(A_{n})-\lambda_{i}(A_{n})|\leq\delta n^{-1/2})=O(n^{o(1)}\delta).$
This result was extended to sparse graphs in [LV20, LL21], and to Wigner matrices (with iid subgaussian entries of mean zero, variance one) in [CJMS24] with better bounds.
From a random matrix perspective, gap sizes were in fact the original motivation for the introduction of random matrices in physics and gap statistics are a staple of the field as can be seen in [Wig55, BS10, AGZ10, EYY12, Tao13, EKYY12] and the references therein. More specifically, the minimal gap in Wigner matrices has also received much attention [Vin01, BAB13, FTW19, Bou22], where in the latest work it was shown to be universal up to some appropriate scaling.
Although these results are quite satisfactory, there remain many interesting open questions, such as to obtain the optimal quadratic repulsion probability bound $O(\delta^{2})$ (under the assumption $\delta\geq n^{{-\omega(1)}}$) in Theorem 1.1, as well as to other models of random matrices with dependent entries. The goal of this paper is to pursue the second direction for a natural model of random matrices that existing results and techniques do not seem to cover – the Laplacian matrix.
Given a graph $G_{n}$ of $n$ vertices, the (combinatorial) Laplacian of $G_{n}$, denoted by $L_{n}$, is the matrix
$L_{n}=D_{n}-A_{n},$
where $D_{n}={\operatorname{diag}}(d_{ii})_{1\leq i\leq n}$ is the diagonal matrix of the vertex degrees, $d_{ii}=\sum_{j=1}^{n}a_{ij}$.
The graph Laplacian is a discrete analogue of the Laplacian operator on manifolds. This matrix is a key object in spectral graph theory, and often of more relevance than the adjacency matrix [Nic18]. The spectrum of this matrix can reveal expansion properties, isoperimetry bounds, random walk mixing times and even the number of spanning trees of graphs [GR01]. Spectral algorithms involving the Laplacian have been spectacularly successful in fundamental tasks such as graph partitioning [HRMP95], clustering [HK92], graph drawing [FPST94], parallel computation [Sim91], graph synchronization [Ban18] and numerous others [MP93]. Laplacians arise in diverse fields ranging from physics to topological data analysis [Fyo99, WW23, BW04].
For the Laplacian operator on manifolds, the eigenvalues and properties of eigenfunctions have been heavily studied in the literature, see for instance [Uhl76, CZY16, Cha84, Zho20, Wan22] and the references therein. One of the most basic questions concerns whether the eigenvalues are separate, and whether the eigenfunctions are Morse functions. Roughly speaking, it has been shown that for a generic metric, the eigenvalues are simple, and the eigenfunctions are smooth. Along these lines, genericity of simple eigenvalues for a metric graph was shown in [Fri05]. Simplicity of eigenvalues and non-vanishing of eigenfunctions of a quantum graph is considered in [BL17].
For (actual) graphs, the “spectral gap" of the Laplacian matrix is the difference between the second smallest eigenvalue and the smallest eigenvalue (which is necessarily zero). This important quantity is tied to the expansion property of the graph and is of interest in many combinatorial optimization problems [Lub12]. However, there are many settings in which the other gaps between consecutive eigenvalues play an important role. In general, the mininmum gap size is intimately tied to the numerical stability of many spectral algorithms [GVL13, Dem97]. For the Laplacian in particular, the minimum gap size or the $k$-th gap size is a crucial parameter or has a specific meaning in various applications. The gap $\lambda_{n-k}-\lambda_{n-k+1}$ indicates the stability of partitioning the graph into $k$ clusters [AEGL21]. Unlike classical random walks, the mixing time for some models of quantum random walks on graphs are governed by the minimal eigenvalue gap of the Laplacian or adjacency matrix [AAKV01, Kem03, CLR20b]. These quantum walks are the basis of efficient quantum algorithms that far outperform classical algorithms and many of these quantum algorithms implicitly assume that the spectrum of the Laplacian matrix for the graphs are simple [AAKV01, CLR20a, CCD+03, Chi04]. For many graph-based learning tasks, graph neural networks (GNNs) are the industry standard solution method [ZCH+20]. Graphs with many symmetries, however, often lead to slow convergence rates and low accuracy [LL22]. One promising solution to this is to endow the vertices with more features to break the symmetries, often by embedding the graph in Euclidean space using the Laplacian matrix [ACGL20, LWWL20]. In this algorithm, the rate of convergence is specified by the size of the smallest eigenvalue gap of the Laplacian matrix [WYZL22, ZL20, MWW23, GBR20]. Some further applications of the minimal eigenvalue gaps are found in graph synchronization [Nyb14, NGB15].
Motivated by all these directions, we show the following
Theorem 1.2 (Eigenvalue gaps of Laplacian, main result).
Let $10<p<$ 1 be fixed and let $A$ be any constant. Let $G_{n}=G(n,p)$ be a random Erdős-Rényi graph with parameter $p$, and let $L_{n}$ be the Laplacian of $G_{n}$. Then for any $\delta\geq n^{-A}$ we have
$\max_{1\leq i\leq n-1}{\mathbb{P}}(|\lambda_{i+1}(L_{n})-\lambda_{i}(L_{n})|\leq\delta n^{-1/2})=O(n^{o(1)}\delta),$
where the implied constants are allowed to depend on $p$ and $A$.
In particular, the spectrum is simple with very high probability. This provides theoretical justification for the empirical success or implicit assumptions in the cited applications. More quantitatively, by taking the union bound over all $i$-th gaps, with probability $1-o(1)$ we see that the minimum gap in fact has order at least $n^{{-3/2}-o(1)}$.
Notations. Throughout this paper, we regard $n$ as an asymptotic parameter going to infinity (in particular, we will implicitly assume that $n$ is larger than any fixed constant, as our claims are all trivial for fixed $n$), and allow all mathematical objects in the paper to implicitly depend on $n$ unless they are explicitly declared to be “fixed” or “constant”. We write $X=O(Y),X\ll Y$, or $Y=\Omega(X),Y\gg X$ to denote the claim that $|X|\leq CY$ for some fixed constant $C$. We write $X=\Theta(Y)$ if $X\gg Y$ and $Y\gg X$. We also use $o(1)$ (and $\omega(1)$ respectively) to denote positive quantities that tend to zero (infinity) as $n\to\infty$.
For a square matrix $X$ of size $n$ and a number $\lambda$, for short we will write $X-\lambda$ instead of $X-\lambda I_{n}$. We denote ${\mathbf{r}}_{i}(X)$ and ${\mathbf{c}}_{i}(X)$ by the $i$-th row and column of $X$. For historical reason, for an $n\times k$ or $k\times n$ matrix $X$ with $k\leq n$, we let $0\leq\sigma_{k}(X)\leq\cdots\leq\sigma_{1}(X)$ be the singular values of $X$, while reversely, for an $n\times n$ symmetric matrix $X$, we let $\lambda_{1}(X)\leq\cdots\leq\lambda_{n}(X)$ denote its eigenvalues and $\delta_{i}(X)=\lambda_{i+1}(X)-\lambda_{i}(X)$ be its $i^{\text{th}}$ eigenvalue gap.
We denote the discrete interval $[m]=\left\{1,2,\cdots,m\right\}$. For ${\mathbf{v}}\in{\mathbb{R}}^{n}$ and $I\subset[n]$, we denote ${\mathbf{v}}_{I}\in{\mathbb{R}}^{I}$ the vector with entries $v_{i},\text{ for }i\in I$. Finally, all the norms in this note, if not specified, will be the usual $l_{2}$-norm.
2. Other main results and our approach
2.1. Eigenvalue gaps for the centered model
It is also natural to study the centered Laplacian (which may go by other names such as Laplacian matrix or Laplacian-type matrix)
$\overline{L}_{n}:=L_{n}-{\mathbf{E}}L_{n}.$
In a more general setting, we can start from a Wigner matrix $X_{n}=(x_{ij})_{1\leq i,j\leq n}$ (with iid entries $x_{ij},1\leq i<j\leq n$ of mean zero and variance one) and form a centered Laplacian $\overline{X}_{n}$ from it by letting $\overline{x}_{ii}=\sum_{j,j\neq i}x_{ij}$ and $\overline{x}_{ij}=-x_{ij}$. Hence our matrix $\overline{L}_{n}/\sqrt{p(1-p)}$ belongs to this family of random matrices.
It is known (see [DJ10]) that the empirical spectral distribution of a random centered Laplacian converges weakly to the free convolution of the semicircular law and the standard gaussian, $\rho_{sc}\boxplus{\mathbf{N}}(0,1)$. The gaps between extreme eigenvalues of the Laplacian have been studied [Jia12, CO07, CLO+22]. Furthermore, among other things, it was shown in [HL20] that individual eigenvalue gaps in the bulk of the spectrum converge to the gap behavior of the GOE. However, these asymptotic results do not cover the basic question of whether $\overline{L}_{n}$ has simple spectrum with high probability. Here we show
Theorem 2.1.
Theorem 1.2 holds for the centered Laplacian model $\overline{L}_{n}$.
We next outline an approach  [^1] There are more advanced approaches to this problem, see for instance [TV11, Eqn. 36]. from [NTV17] to prove Theorem 1.1, and single out the difficulty of the current model. Assume that $X_{n}$ is a symmetric/Hermitian matrix where $\lambda_{i+1}-\lambda_{i}\leq\delta/n^{1/2}$.
Write
$X_{n}=\begin{pmatrix}X_{n-1}&{\mathbf{c}}_{n}\\ {\mathbf{c}}_{n}^{*}&x_{nn}\end{pmatrix}$ (1)
where ${\mathbf{c}}_{n}$ is the last column vector. From the Cauchy interlacing law, we observe that $\lambda_{i}(X_{n})\leq\lambda_{i}(X_{n-1})\leq\lambda_{i+1}(X_{n})$. Let ${\mathbf{u}}$ be the (unit) eigenvector of $\lambda_{i}(X_{n})$; we write ${\mathbf{u}}=({\mathbf{w}},b)$, where ${\mathbf{w}}$ is a vector of length $n-1$ and $b$ is a scalar. We have
$\begin{pmatrix}X_{n-1}&{\mathbf{c}}_{n}\\ {\mathbf{c}}_{n}*&x_{nn}\end{pmatrix}\begin{pmatrix}{\mathbf{w}}\\ b\end{pmatrix}=\lambda_{i}(X_{n})\begin{pmatrix}{\mathbf{w}}\\ b\end{pmatrix}.$
Extracting the top $n-1$ components of this equation we obtain
$(X_{n-1}-\lambda_{i}(X_{n})){\mathbf{w}}+b{\mathbf{c}}_{n}=0.$
Let ${\mathbf{v}}$ be the unit eigenvector of $X_{n-1}$ corresponding to $\lambda_{i}(X_{n-1})$. By multiplying with ${\mathbf{v}}^{T}$, we obtain
$|b{\mathbf{v}}^{T}{\mathbf{c}}_{n}|=|{\mathbf{v}}^{T}(X_{n-1}-\lambda_{i}(X_{n})){\mathbf{w}}|=|\lambda_{i}(X_{n-1})-\lambda_{i}(X_{n})||{\mathbf{v}}^{T}{\mathbf{w}}|.$ (2)
We conclude that, if $\lambda_{i+1}-\lambda_{i}\leq\delta n^{-1/2}$ holds, then $|b{\mathbf{v}}^{T}{\mathbf{c}}_{n}|\leq\delta n^{-1/2}$. If we assume that $|b|\gg n^{-1/2+o(1)}$, then
$|{\mathbf{v}}^{T}{\mathbf{c}}_{n}|\ll n^{o(1)}\delta.$ (3)
Thus we have reduced the problem to bounding the probability with respect to $X_{n-1}$ and ${\mathbf{c}}_{n}$ that $X_{n-1}$ has an eigenvector ${\mathbf{v}}$ and such that $|{\mathbf{v}}^{T}{\mathbf{c}}_{n}|=O(n^{o(1)}\delta)$. The plan is then divided into two separate steps:
  1. (1)
    Step 1. With extremely high probability with respect to $X_{n-1}$, all of the unit eigenvectors ${\mathbf{v}}$ of $X_{n-1}$ do not have structure (where we will delay the precise definition of this structure to a later discussion);
  2. (2)
    Step 2. Conditioned on $X_{n-1}$ from the first step, the probability with respect to ${\mathbf{c}}_{n}$ that (3) holds is bounded by $O(\delta)$.
The structure of the Laplacian matrix $L_{n}$ (and $\overline{L}_{n}$) pose several challenges when compared with that of the symmetric/Hermitian matrix $A_{n}$ (which plays the role of $X_{n}$ above). The most apparent difficulty is the dependence among the entries above the diagonal. For instance if we fix all the entries of $L_{n-1}$, the upper left principle minor of size $n-1$ of $L_{n}$, then $L_{n}$ is determined. In other words, even if one can execute Step 1 for $L_{n-1}$, we cannot rely on Step 2 because ${\mathbf{c}}_{n}(L_{n})$ is already determined after fixing $L_{n-1}$. To avoid this deadlock, we will proceed as follows (using the model $G(n,1/2)$ as an example – the approach naturally extends to $G(n,p)$ and beyond). First, assume that the vertices of $G=G_{n}$ are ordered as $(v_{1},\dots,v_{n})$. We sample the entries $a_{ij}$ of the adjacency matrix (via $G(n,1/2)$) and compute the degrees $d_{1},\dots,d_{n}$ of the vertices, and then form $\overline{L}_{n}$. Let $\overline{L}_{n}^{\prime}$ be the matrix obtained from $L_{n}$ by fixing all rows and columns, except that the last two columns ${\mathbf{c}}_{n-1},{\mathbf{c}}_{n}$ (as well as the last two rows ${\mathbf{r}}_{n-1},{\mathbf{r}}_{n}$) are replaced by ${\mathbf{c}}_{n-1}+{\mathbf{c}}_{n}$ and ${\mathbf{c}}_{n}-{\mathbf{c}}_{n-1}$  [^2] In fact, to preserve the spectrum, we will replace ${\mathbf{c}}_{n-1},{\mathbf{c}}_{n}$ by $({\mathbf{c}}_{n-1}+{\mathbf{c}}_{n})/\sqrt{2}$ and $({\mathbf{c}}_{n}-{\mathbf{c}}_{n-1})/\sqrt{2}$ (see Section 7), but let’s ignore this minor problem for now as the proofs remain the same..
Next, we reshuffle the neighbors of $v_{n-1}$ and $v_{n}$ as follows: consider the set $I_{n-2}$ of indices $21\leq j\leq n-$ 2 where $v_{j}$ is connected to exactly one of $v_{n-1}$ or $v_{n}$. Then for each $j\in I_{n-2}$, we flip a fair coin to either keep or swap whether each of $(v_{j},v_{n-1})$ and $(v_{j},v_{n})$ is an edge (see Figure 1). In other words, if $X=(x_{1},\dots,x_{n-2})$ and $Y=(y_{1},\dots,y_{n-2})$ are the (restricted to the first $n-2$ coordinates) column vectors of $A_{n}$ (or $L_{n},\overline{L}_{n}$) associated to $v_{n-1}$ and $v_{n}$, then $I_{n-2}$ is the collection of indices $j$ where $(x_{j},y_{j})=(0,1)$ or $(1,0)$ (respectively $(0,-1)$ or $(-1,0)$ in the $L_{n}$ case, and $(1/2,-1/2)$ or $(-1/2,1/2)$ in the $\overline{L}_{n}$ case). We then flip a fair coin to reassign $(x_{j},y_{j})$ to $(0,1)$ or $(1,0)$. To some extent, this reshuffling is similar to [McK81] and [Coo17, LLT+21, NW22, NW24] where shufflings/switchings were used within random graphs and random matrices. However our implementation here is rather straightforward. We can easily see that the above process does not change the law of $\overline{L}_{n}$. In terms of $\overline{L}_{n}^{\prime}$, this process searches for the index set $I_{n-2}\subset[n-2]$ where the entries of ${\mathbf{c}}_{n-1}(\overline{L}^{\prime}_{n})$ (or the entries of ${\mathbf{c}}_{n-1}(\overline{L}_{n})+{\mathbf{c}}_{n}(\overline{L}_{n})$) is zero, and then for each $j\in I_{n-2}$, the entries of ${\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})$ take values ±$\pm 1$ (independently from each other, and from $\overline{L}^{\prime}_{n-1}$) with probability 1/2. By Chernoff’s bound, with high probability $I_{n-2}$ has size approximately $(n-2)/2$ (or $(n-2)p$ in general). With the new randomness of ${\mathbf{c}}_{n}(\overline{L}_{n}^{\prime})_{|_{I_{n-2}}}$, we can now move to Step 2 to study (3), which can be rewritten as
$|{\mathbf{v}}^{T}_{|_{I_{n-2}}}{\mathbf{c}}_{n}(\overline{L}_{n}^{\prime})_{|_{I_{n-2}}}-a|\ll\delta$ (4)
for some deterministic $a$.
While we can proceed with Step 2, a new problem arises, that in Step 1 it is not enough to know that ${\mathbf{v}}$ does not have structure, but we have to show that ${\mathbf{v}}_{I}$ does not have structure for any eigenvector of $\overline{L}_{n-1}^{\prime}$. As far as we are concerned, the property of eigenvectors not having local structure has not been studied before in the literature, except in the no-gaps delocalization aspect. (Although in the finite field setting, this local aspect has been studied in [NW22] for null vectors.)
$I_{n-2}$.$v_{j}$.$v_{n-1}$.$v_{n}$
$I_{n-2}$.$v_{j}$.$v_{n-1}$.$v_{n}$
Figure 1. Reshuffling neighbors.
2.2. No-gaps and no-structure delocalization
Our plan for Step 1 of the modified plan has two substeps
  • Substep 1(i). We first show that with high probability, $\|{\mathbf{v}}_{I}\|_{2}$ is not too small for all $I$ of size $\Theta(n)$ (Theorem 2.3).
  • Substep 1(ii). We then show that ${\mathbf{v}}_{I}$ does not have structure (Theorem 2.4).
The first type of result is called no-gaps delocalization, a notion pioneered by Rudelson and Vershynin in [RV16]. Note that in Random Matrix Theory, delocalization usually means that the eigenvectors are flat, not localized. For instance for the current centered Laplacian model it has been shown in [HL20] that all unit eigenvectors ${\mathbf{v}}$ (corresponding to the eigenvalues in the bulk) are completely delocalized in the sense that $\|{\mathbf{v}}\|_{\infty}\leq n^{{-1/2+o(1)}}$ with very high probability. So $\|{\mathbf{v}}_{I}\|_{2}$ is small over any set $I$ of size $o(n)$. The no-gaps delocalization, on the other hand, addresses the property that $\|{\mathbf{v}}_{I}\|_{2}$ cannot be small over any set $I$ of order $n$. We will cite below a very strong result of Rudelson-Vershynin in [RV16], that works for any general random non-symmetric, symmetric and skew-symmetric matrices $X_{n}$ (where $x_{ij}$ and $x_{ji}$ can depend on each other, but otherwise are independent, subgaussian of mean zero and variance one).
Theorem 2.2.
There exist constants $C,C^{\prime},c$ such that the following holds. Let $\lambda_{n}\geq\frac{1}{n}$ and $s\geq C\lambda_{n}^{-7/6}n^{-1/6}+\exp(-c/\sqrt{\lambda_{n}})$. Then with probability at least $1-(C^{\prime}s)^{\lambda_{n}n}$, every unit eigenvector ${\mathbf{v}}$ of $X_{n}$
$\|{\mathbf{v}}_{I}\|_{2}\geq(\lambda_{n}s)^{6}$
for all $I\subset[n]$ such that $|I|\geq\lambda_{n}n$.
While this result is extremely strong, the method of [RV16] does not seem to extend to (i) matrices where the diagonal entries depend on other entries (the way $\overline{L}_{n}$ does), and (ii) matrices of norm of order $\omega(\sqrt{n})$ – while here with high probability $\overline{L}_{n}$ has norm of order $\Theta(\sqrt{n\log n})$.
In this note we show the following analogue
Theorem 2.3 (Affine no-gaps delocalization of eigenvectors of $\overline{L}^{\prime}_{n-1}$).
Let $n_{0}=\lambda_{n}n$ where
$\frac{1}{\log\log n}\leq\lambda_{n}<1.$
The probability with respect to $\overline{L}^{\prime}_{n-1}$ (or $\overline{L}_{n-1}$) that it has a unit eigenvector ${\mathbf{v}}$ and an index set $I\subset[n]$ of size $n_{0}$ such that ${\mathbf{v}}\perp\mathbf{1}$ and
$\inf_{a\in{\mathbb{R}}}\|{\mathbf{v}}_{I}-a1_{I}\|_{2}\leq\frac{1}{(\log n)^{\frac{A_{0}}{\lambda_{n}}}}$ (5)
for some sufficiently large constant $A_{0}$, is bounded by $n^{-\omega(1)}$ as $n\to\infty$.
As a corollary, when $\lambda_{n}$ is of constant order, with high probability for any non-trivial unit eigenvector ${\mathbf{v}}$, not only is the total mass over $I$ not small, $\|{\mathbf{v}}_{I}\|_{2}\geq\frac{1}{(\log n)^{O(1)}}$, but it is also not small after shifting (affine delocalization)
$\inf_{a\in{\mathbb{R}}}\|{\mathbf{v}}_{I}-a1_{I}\|_{2}\geq\frac{1}{(\log n)^{O(1)}}.$
This translates into a statement about “incompressibility” of affine shifts of segments of eigenvectors. It is possible that the result continues to hold for $\inf_{a\in{\mathbb{R}}}\|{\mathbf{v}}_{I}-a1_{I}\|_{2}\geq\Theta_{\lambda_{n}}(1)$ and for $\lambda_{n}\ll\frac{1}{\log\log n}$ (with possibly much finer approximation than (5)) as in [RV16] but it seems $\frac{1}{(\log n)^{O(1)}}$ (for the mass contribution) and $\frac{1}{\log\log n}$ (for the lower bound of $\lambda_{n}$) are the limit of our proof. Theorem 2.3 applies to subsets $I$ of size larger than $n/\log\log n$ which is enough for our purposes, and also, the affine aspect is new and will play a key role for later parts of the paper. While the result is much weaker compared to Theorem 2.2, our matrix model, the random Laplacian, seems to be more involved and we will have to develop new ways to handle these difficulties. Furthermore, the behavior of eigenvectors for the Laplacian are fundamentally different than that of the non-symmetric or Wigner matrices. Even for a Laplacian matrix generated from gaussian random variables, the eigenvectors are not uniformly distributed over the sphere. In fact, simulations show that eigenvectors corresponding to large eigenvalues appear to be mildly localized.
Next, we will discuss Substep 1(ii), which is another innovative aspect of our note. To do this, we first introduce an important concept following Rudelson and Vershynin.
Definition 2.3.
For a unit vector ${\mathbf{x}}\in{\mathbb{R}}^{n}$, we define the Least Common Denominator of ${\mathbf{x}}$ with parameters $\kappa>0$ and $\gamma\in(0,1)$ as
$\mathbf{LCD}_{\kappa,\gamma}({\mathbf{x}})=\inf\left\{\theta>0:{\operatorname{dist}}(\theta{\mathbf{x}},{\mathbb{Z}}^{n})<\min\{\gamma\|\theta{\mathbf{x}}\|_{2},\kappa\}\right\}$
In this paper, if not specified otherwise, we will choose $\kappa=\kappa_{n}=n^{c}$ for some sufficiently small constant $c$, while $\gamma=\gamma_{n}$ is sufficiently small, which can tend to zero slowly.
One of the highlights of the paper is to show that not only are the eigenvectors no-gaps delocalized, but also “no-structure delocalized”.
Theorem 2.4 (No-structure delocalization of Laplacian eigenvectors).
Let $A>0$ be given. Let $n_{0}=\lambda_{n}n$ where
$\frac{1}{\sqrt{\log\log n}}\leq\lambda_{n}<1.$
The probability with respect to $\overline{L}^{\prime}_{n-1}$ (or $\overline{L}_{n-1}$) that it has a unit eigenvector ${\mathbf{v}}$ and an index set $I\subset[n]$ of size $n_{0}$ such that ${\mathbf{v}}\perp\mathbf{1}$ and
$\mathbf{LCD}_{\kappa,\gamma}({\mathbf{v}}_{I})\geq n^{A}$
is bounded by $n^{-\omega(1)}$ as $n\to\infty$.
Here the bound $\frac{1}{\sqrt{\log\log n}}\leq\lambda_{n}$ is likely to be improvable to $\frac{1}{\log\log n}\leq\lambda_{n}$ as in Theorem 2.3 but we will not focus on this.
2.4. Approximated eigenvetors and more general results
Given $A$ from Theorems 2.3 and 2.4, we choose $C$ to be a sufficiently large constant. By approximating the eigenvalues $\lambda$ with scale $n^{-C}$ and using Lemma 3.2, it suffices to consider the event that there exists ${\mathbf{v}}\in S^{n-2}$ and $\lambda_{0}\in{\mathcal{N}}_{\lambda}=\{\frac{k}{n^{C}},k\in{\mathbb{Z}},|k|\leq n^{C+1}\}$ such that $|(\overline{L}^{\prime}_{n-1}-\lambda_{0}){\mathbf{v}}|\leq n^{-C}$, and $v_{1}+\dots+v_{n}=0$  [^3] As the size of ${\mathcal{N}}_{\lambda}$ is of order $|{\mathcal{N}}_{\lambda}|=n^{O(1)}$, while our probability is of order $n^{-\omega(1)}$, it suffices to fix one $\lambda_{0}$ from ${\mathcal{N}}_{\lambda}$.. Let
$D:=\lceil n^{C+1/2}\rceil.$ (6)
By further $\sqrt{n}/D$-approximating ${\mathbf{v}}$ in $l_{2}$-norm, we can assume that the entries of ${\mathbf{v}}$ are in $\frac{1}{D}{\mathbb{Z}}$, and
$\|(\overline{L}^{\prime}_{n-1}-\lambda_{0}){\mathbf{v}}\|_{2}=O(\sqrt{n}\log n\times n^{-C});\mbox{ and }|\sum_{i}v_{i}|\leq n/D=n^{1/2-C};\mbox{ and }\|{\mathbf{v}}\|_{2}=1+O(n^{-C}).$ (7)
We call the above ${\mathbf{v}}$ approximated non-trivial eigenvectors.
Theorem 2.5.
Theorems 2.3 and 2.4 hold for approximated eigenvectors from (7) of $\overline{L}^{\prime}_{n-1}$ (or of $\overline{L}_{n-1}$).
2.5. Overcrowding estimate
One of the main obstacles to rule out the event of having $\|{\mathbf{v}}_{I}\|_{2}=o(1)$ or $\mathbf{LCD}({\mathbf{v}}_{I})$ small is that we have no information on the remaining vector ${\mathbf{v}}_{I^{c}}$. For the sake of discussion, let us assume that $X_{n}{\mathbf{v}}=0$. We can rewrite this as $M_{11}{\mathbf{v}}_{I}+M_{21}{\mathbf{v}}_{I^{c}}=0$ and $M_{12}{\mathbf{v}}_{I}+M_{22}{\mathbf{v}}_{I^{c}}=0$, where
[…]
Given ${\mathbf{v}}_{I}$ from some finite set for which we can control the size, conditioning on $M_{21},M_{12}$ and $M_{22}$ we can solve for ${\mathbf{v}}_{I^{c}}$ from $M_{21}{\mathbf{v}}_{I}+M_{22}{\mathbf{v}}_{I^{c}}=0$, that ${\mathbf{v}}_{I^{c}}=-M_{22}^{{-1}}M_{21}{\mathbf{v}}_{I}$. Substitute this into the first equation, we hence obtain
$(M_{11}-M_{12}M_{22}^{{-1}}M_{21}){\mathbf{v}}_{I}=0.$
We can then bound this probability using the randomness on the off-diagonal entries of $M_{11}$.
One of the key problems of this approach, taking into account that we are working with asymptotic approximations rather than with identities, is that we have to control the norm of $\|M_{22}^{-1}\|_{2}$. This seems to be a very hard problem, where it is only known recently from [NW24] that $M_{22}$ is invertible with high probability. To avoid this difficulty, we will not work with $M_{22}$ directly, but a near-square rectangular submatrix $M_{22}^{\prime}$ of it, where we can control the non-zero least singular values quite effectively. One drawback of this modification is that we can no longer solve for ${\mathbf{v}}_{I^{c}}$, but for only (and approximately) a major part of it. However, this problem can be resolved by again passing to another approximated eigenvector. We now state the main result on the strong invertibility of $M_{22}^{\prime}$ in the form of $L_{n}$.
Theorem 2.6 (Overcrowding for spectrum of Laplacian matrices).
Let $L_{n}$ be the Laplacian corresponding an Erdős-Rényi graph $G(n,p)$. There exist constants $C\geq 1$, $c>0$, depending only on $p$ such that
$\mathbb{P}\left(\sigma_{n-k+1}(L_{n}+F)\leq\frac{ck}{\sqrt{n}}\right)=O\Big{(}\exp(-\Theta(k^{3/2}/\log n)\Big{)},$
for $k\geq C\log n$ and where the implied constants depend on $C,c$.
This result, which can be seen as an analogue of Theorem 1.7 in [Ngu18] where iid and symmetric matrices were considered, is of its own interest. Interestingly, as we can see, the result works for all choices of deterministic matrices $F$, and hence works for $\overline{L}_{n}$ as well  [^4] The proof automatically extends to the model $\overline{L}^{\prime}_{n}$..
The dependence on $k$ in the exponential probability bound in Theorem 2.6 can likely be improved to $k^{2}$ as in [Ngu18]. However, we do not pursue this extension further in this paper as for our application, it suffices that when $k=\Theta(n/\log^{C}n)$, the probability bound is super exponentially small.
2.6. Small coordinates of eigenvectors
As first observed in [NTV17], a direct consequence of controlling the structure of the eigenvectors of a principle minor is that the eignevectors of the Laplacian cannot have many coordinates of small size (in fact, we show they cannot have more than one zero coordinate). The zero-coordinates of Laplacian eigenvectors play a special role in nodal domains in spectral graph theory and the physics of mechanical systems, where they are referred to as soft nodes [DLL11, BykgHL+04, CKK19, CK23]. For dynamical systems on graphs governed by the Laplacian, soft nodes are of interest theoretically and practically because they are immune to forcing and damping at that node [CKS12].
In the random matrix setting, the following result can be seen as another instance of no-gaps delocalization in the extreme case where the subset is of constant size.
Theorem 2.7 (Small coordinates of eigenvectors).
For any $A$, there exists a $B$, depending on $A$, such that with probability at least $1-n^{-A}$, an eigenvector of $L_{n}$ or $\overline{L}_{n}$ does not have more than one coordinate of size less than $n^{-B}$.
Naturally, we expect that the reality is that there are no small coordinates with high probability, as was shown for Wigner matrices in [NTV17].
2.7. Further Remarks
By following our method, the no-structure delocalization result, Theorem 2.4, should hold for other models such as non-symmetric, symmetric, and skew symmetric matrices of independent subgaussian entries of mean zero and variance one. In fact, we suspect that for these models, the result should be stronger, that $\lambda_{n}$ can be as small as $n^{-c}$ for some positive constant $c$ (or even as small as $n^{-1+c}$), and the probability bound $n^{-\omega(1)}$ might be improved to a subexponential rate at least.
Additionally, it seems to be an important problem to extend these results to Laplacians of sparse graphs $G(n,p)$, where $p\rightarrow 0$ as $n\rightarrow\infty$, as well as to random $d$-regular graphs (for either fixed $d$ or $d\to\infty$ with $n$). Finally, it is of interest to study the law of the minimum gap of these random Laplacians, with a focus on demonstrating universality.
3. Supporting lemmas
We will frequently make use of the following deterministic lemma.
Lemma 3.1 (Theorem 6 of [NY17]).
Let $A:{\mathbb{R}}^{n}\to{\mathbb{R}}^{k}$ of full rank $k$. Then for every $11\leq l\leq k-$ 1, there exists a $I\subseteq\{1,\ldots,n\}$ of size $l$ such that if $A_{I}$ denotes $A\mid_{{\mathbb{R}}^{I}}$, then
$\sigma_{l}(A_{I})\geq c\max_{r\in\{l+1,\ldots,k\}}\sqrt{\frac{(r-l)\sum_{i=r}^{n}\sigma_{i}(A)^{2}}{nr}}$
for an absolute constant $c$.
Next we need the following elementary fact for the centered Laplacian matrices.
Lemma 3.2 (Norm of centered Laplacian matrix).
We have
${\mathbb{P}}(\|\overline{L}_{n}\|_{2}\geq\lambda\sqrt{n}\log n)\ll\exp(-\Theta(\lambda^{2}\log n)).$
In particular, with probability at least $1-\exp(-\Theta(\log^{2}n))$, all submatrices of $\overline{L}_{n}$ have norm, say, bounded by $\sqrt{n}\log n$.
Proof.
(of Lemma 3.2) By the triangle inequality
$\|\overline{L}_{n}\|_{2}\leq\|D_{n}-{\mathbf{E}}D_{n}\|_{2}+\|A_{n}-{\mathbf{E}}A_{n}\|_{2}.$
We can then bound $\|D_{n}-{\mathbf{E}}D_{n}\|_{2}$ and $\|A_{n}-{\mathbf{E}}A_{n}\|_{2}$ separately. Alternatively, it follows from [Ban18, Theorem 3.2] that the median of $\|\overline{L}_{n}\|_{2}$ has order $\Theta(\sqrt{n\log n})$. Then by Talagrand concentration theorem, as $\|\overline{L}_{n}\|_{2}$ is convex and $\sqrt{n}$-Lipschitz with respect to the .$\|.\|_{HS}$-norm, we have
${\mathbb{P}}(|\|\overline{L}_{n}\|_{2}-M(\|\overline{L}_{n}\|_{2})|\geq\sqrt{n}\lambda)\leq\exp(-\Theta(\lambda^{2})).$
Occasionally, in many proofs we will need the following tensorization Lemma to transform anti-concentration bound from independent random variables to random vectors ([RV08, Lemma 3.4]).
Lemma 3.3.
Let $X=(x_{1},\ldots,x_{n})$ be a random vector in ${\mathbb{R}}^{n}$ with independent coordinates $X_{k}$. Assume that ${\mathbb{P}}(|x_{i}|<\delta)\leq K\delta$ for all $\delta\geq\delta_{0}$. Then
${\mathbb{P}}(x_{1}^{2}+\dots+x_{n}^{2}<\delta^{2}n)\leq(C_{0}K\delta)^{n}$
for some absolute constant $C_{0}$.
Proof.
(of Lemma 3.3) Let $\delta\geq\delta_{0}$, we have the following by Chebychev’s inequality:
${\mathbb{P}}\Big{(}\sum_{i=1}^{n}x_{i}^{2}<\delta^{2}n\Big{)}={\mathbb{P}}\Big{(}n-\frac{1}{\delta^{2}}\sum_{i=1}^{n}x_{i}^{2}>0\Big{)}\leq\mathbb{E}\exp\Big{(}n-\frac{1}{\delta^{2}}\sum_{i=1}^{n}x_{i}^{2}\Big{)}=e^{n}\prod_{i=1}^{n}\mathbb{E}\exp(-x_{i}^{2}/\delta^{2}).$
Integrating the tail bound we have
[…]
Decomposing the integration and using the assumption that ${\mathbb{P}}(|x_{i}|<\delta u)\leq{\mathbb{P}}(|x_{i}|<\delta)\leq K\delta$ for $u\in[0,1]$, we have
[…]
Plugging this back to the first inequality, we have
${\mathbb{P}}\Big{(}\sum_{i=1}^{n}x_{i}^{2}<\delta^{2}n\Big{)}<e^{n}(C_{0}K\delta)^{n}.$
For the remainder of this section we will discuss several properties of vectors having small $\mathbf{LCD}$, including compressible and incompressible vectors. First, the following vectors are outputs of our no-gaps delocalization result Theorem 2.3.
Claim 3.4.
Let $0<c_{1}\leq c_{0}<1<C$ be given parameters (that might depend on $n$). Assume that ${\mathbf{x}}\in S^{n-1}$ be such that $\|{\mathbf{x}}_{I}\|_{2}\geq c_{1}$ for any $I\subset[n]$ of size at least $\lfloor c_{0}n\rfloor$. Then the following holds.
  • For any set $S$ of size $\lfloor c_{0}n\rfloor$, there exists one $i\in S$ such that $|x_{i}|\geq(c_{1}/c_{0})^{{1/2}}/\sqrt{n}$. It thus follows that for all but $\lfloor c_{0}n\rfloor-1$ indices $i$ we have $|x_{i}|\geq(c_{1}/c_{0})^{{1/2}}/\sqrt{n}$.
  • The set $S_{c_{0},c_{1},C}$ of indices $i$ where
    $(c_{1}/c_{0})^{{1/2}}/\sqrt{n}\leq|x_{i}|\leq C/\sqrt{n}$
    has size at least $(1-c_{0}-C^{-2})n$.
Proof.
It suffices to show the first part. Assume otherwise, then we would have $\|{\mathbf{x}}_{S}\|^{2}\leq|S|(c_{1}/c_{0})/n<c_{1}$, contradiction. ∎
Given parameters $c_{0},c_{1},C$, we define the $\mathrm{spread}_{c_{0},c_{1},C}({\mathbf{x}})$ to be the set of coordinates $k$ satisfying
$(c_{1}/c_{0})^{{1/2}}/\sqrt{n}\leq|x_{i}|\leq C/\sqrt{n}.$
The next Lemma connects the notion of compressible vectors and $\mathbf{LCD}$ together.
Lemma 3.5.
Assume that ${\mathbf{x}}$ satisfies Claim 3.4. Then
$\mathbf{LCD}_{\kappa,\gamma}({\mathbf{x}})\geq\frac{\sqrt{n}}{2C}$
for any $\kappa>0$ and $\gamma\leq\frac{1}{2}[(1-c_{0}-C^{-2})(c_{1}/c_{0})]^{{1/2}}$.
It is worth noting that the smaller $\gamma$ is, the larger the $\mathbf{LCD}_{\kappa,\gamma}$ becomes.
Proof.
(of Lemma 3.5) For each $k\in\mathrm{spread}({\mathbf{x}})$,
$(c_{1}/c_{0})^{{1/2}}/\sqrt{n}\leq|x_{k}|\leq C/\sqrt{n}.$
For all $\theta\in(0,\frac{\sqrt{n}}{2C})$, and $k\in\mathrm{spread}({\mathbf{x}})$,
$\mathrm{dist}(\theta x_{k},{\mathbb{Z}})=|\theta x_{k}|\geq\theta(c_{1}/c_{0})^{{1/2}}/\sqrt{n}.$
Therefore, for $\theta\in(0,\frac{\sqrt{n}}{2C})$,
$\mathrm{dist}(\theta{\mathbf{x}},{\mathbb{Z}}^{n})\geq\mathrm{dist}(\theta{\mathbf{x}}|_{\mathrm{spread}({\mathbf{x}})},{\mathbb{Z}}^{\mathrm{spread}({\mathbf{x}})})\geq(1-c_{0}-C^{-2})^{1/2}(c_{1}/c_{0})^{{1/2}}\theta>\gamma\theta.$
Therefore,
$\mathbf{LCD}_{\kappa,\gamma}({\mathbf{x}})\geq\frac{\sqrt{n}}{2C}.$
Fact 3.6 ($\mathbf{LCD}$ of subvectors).
Assume that ${\mathbf{x}}^{\prime}$ is a subvector of ${\mathbf{x}}$ such that $\gamma\|{\mathbf{x}}\|_{2}/\|{\mathbf{x}}^{\prime}\|_{2}<1$. We have
$\mathbf{LCD}_{\kappa,\gamma\|{\mathbf{x}}\|_{2}/\|{\mathbf{x}}^{\prime}\|_{2}}(\frac{{\mathbf{x}}^{\prime}}{\|{\mathbf{x}}^{\prime}\|_{2}})\leq\frac{\|{\mathbf{x}}^{\prime}\|_{2}}{\|{\mathbf{x}}\|_{2}}\mathbf{LCD}_{\kappa,\gamma}(\frac{{\mathbf{x}}}{\|{\mathbf{x}}\|_{2}}).$
Proof.
(of Fact 3.6) It suffices to assume $\|{\mathbf{x}}\|_{2}=1$. Note that for any $D>0$, if ${\operatorname{dist}}(D{\mathbf{x}},{\mathbb{Z}}^{n})\leq\min(\gamma D,\kappa)$ then trivially
${\operatorname{dist}}((D\|{\mathbf{x}}^{\prime}\|_{2})({\mathbf{x}}^{\prime}/\|{\mathbf{x}}^{\prime}\|_{2}),{\mathbb{Z}}^{n})\leq\min((\gamma/\|{\mathbf{x}}^{\prime}_{2}\|_{2})\|(D\|{\mathbf{x}}^{\prime}\|_{2})({\mathbf{x}}^{\prime}/\|{\mathbf{x}}^{\prime}\|_{2})\|_{2},\kappa).$
One of the main features of $\mathbf{LCD}$ is that it detects the level of radius where we can bound the small ball probability asymptotically optimally (up to a multiplicative constant).
Theorem 3.7 (Small ball probability via LCD, [RV08]).
Let $\xi$ be a sub-gaussian random variable of mean zero and variance one, and let $\xi_{1},\cdots,\xi_{n}$ be iid copies of $\xi$. Consider a vector ${\mathbf{x}}\in{\mathbb{R}}^{n}$ which satisfies $\left\|{\mathbf{x}}\right\|\geq 1$. Then, for every $\kappa>0$ and $\gamma\in(0,1)$, and for
$\varepsilon\geq\frac{1}{\mathbf{LCD}_{\kappa,\gamma}({\mathbf{x}})},$
we have
$\rho_{\varepsilon}({\mathbf{x}}):=\mathcal{L}(\sum_{i=1}^{n}x_{i}\xi_{i},\varepsilon):=\sup_{a\in{\mathbb{R}}}\mathbb{P}\left(\left|\sum_{i=1}^{n}x_{i}\xi_{i}-a\right|\leq\varepsilon\right)=O\left(\frac{\varepsilon}{\gamma}+e^{-\Theta(\kappa^{2})}\right),$
where the implied constants depend on $\xi$.
Finally, we need the following important results, whose proof will be given for the reader’s convenience.
Lemma 3.8 (Nets of structured vectors).
Let $\gamma,\alpha,\kappa,D_{0}$ be given positive parameters, where $\gamma<1$, and $\kappa$ and $D_{0}$ are sufficiently large. The following holds for $m$ sufficiently large.
  • (Nets of vectors having $\mathbf{LCD}$ belonging to $(D_{0},2D_{0}]$, [RV09, Lemma 4.7][NTV17, Lemma 11.3]) Assume that $\kappa/D_{0}\leq 1/2$. Let $S_{\alpha,D_{0}}$ be the set of ${\mathbf{x}}\in{\mathbb{R}}^{m}$ with $\|{\mathbf{x}}\|=\alpha$ and that
    $\mathbf{LCD}_{\kappa,\gamma}({\mathbf{x}}/\|{\mathbf{x}}\|_{2})\in[D_{0},2D_{0}].$
    Then there exists a $(2\alpha\kappa/D_{0})$-net of $S_{D_{0}}$ of cardinality at most $(C_{0}D_{0}/\sqrt{m})^{m}$, where $C_{0}$ is an absolute constant. Consequently, there exists a $(2\kappa/D_{0})$-net of $S_{\alpha,D_{0}}$ of cardinality at most $(C_{0}(\alpha+1)D_{0}/\sqrt{m})^{m}$.
  • (Nets of unit vectors having $\mathbf{LCD}$ smaller than $D$, [NTV17, Lemma 11.4]) Assume that $2\kappa\leq D_{0}\leq D$. Then the set $S_{1,D_{0}}$ has a $(2\kappa/D)$-net of cardinality at most $(C_{0}D/\sqrt{m})^{m}$ for some absolute constant $C_{0}$.
  • (Trivial nets) The number of vectors in $(\frac{1}{D}{\mathbb{Z}})^{m}$ that have norm bounded by $\alpha$ is at most $(C^{\prime}\alpha D/\sqrt{m}+1)^{m}$, and hence this set of vectors accepts an $\sqrt{m}/D$-net of size $(C^{\prime}\alpha D/\sqrt{m}+1)^{m}$.
Proof.
(of Lemma 3.8) Let us first focus on the net of $S_{\alpha,D_{0}}$. For ${\mathbf{x}}\in S_{\alpha,D_{0}}$ and $\|{\mathbf{x}}\|_{2}=\alpha$, denote
$D({\mathbf{x}}):=\mathbf{LCD}_{\kappa,\gamma}({\mathbf{x}}/\alpha).$
By definition, $D_{0}\leq D({\mathbf{x}})\leq 2D_{0}$ and there exists $p\in{\mathbb{Z}}^{m}$ with
$\left\|\frac{{\mathbf{x}}}{\alpha}-\frac{p}{D({\mathbf{x}})}\right\|_{2}\leq\frac{\kappa}{D({\mathbf{x}})}.$
This implies that $\|p\|\approx D({\mathbf{x}})$, more precisely
$1-\frac{\kappa}{D({\mathbf{x}})}\leq\left\|\frac{p}{D({\mathbf{x}})}\right\|_{2}\leq 1+\frac{\kappa}{D({\mathbf{x}})}.$ (8)
This implies that
$\|p\|_{2}\leq 3D({\mathbf{x}})/2\leq 3D_{0}.$ (9)
It also follows from (8) that
$\left\|{\mathbf{x}}-\alpha\frac{p}{\|p\|}\right\|_{2}\leq\alpha\left\|\frac{{\mathbf{x}}}{\alpha}-\frac{p}{D({\mathbf{x}})}\right\|_{2}+\alpha\left\|\frac{p}{\|p\|}(\frac{\|p\|}{D({\mathbf{x}})}-1)\right\|_{2}\leq\frac{2\alpha\kappa}{D({\mathbf{x}})}\leq\frac{2\alpha\kappa}{D_{0}}.$ (10)
Now set
${\mathcal{N}}_{0}:=\left\{\alpha\frac{p}{\|p\|},p\in{\mathbb{Z}}^{m}\cap B(0,3D_{0})\right\}.$
By (9) and (10), ${\mathcal{N}}_{0}$ is a $\frac{2\alpha\kappa}{D_{0}}$-net for $S_{D_{0}}$. On the other hand, it is known that the size of ${\mathcal{N}}_{0}$ is bounded by $(C_{0}\frac{D_{0}}{\sqrt{m}})^{m}$ for some absolute constant $C_{0}$.
For the consequence, it suffices to assume $\alpha\geq 1$. As we can cover $S_{\alpha,D_{0}}$ by $(C_{0}D_{0}/\sqrt{m})^{m}$ balls of radius $2\alpha\kappa/D_{0}$, we can then cover these balls by smaller balls of radius $2\kappa/D$, the number of such small balls is at most $(O(\alpha))^{m}$. Thus there are at most $(C_{0}\frac{\alpha D_{0}}{\sqrt{m}})^{m}$ balls of radius $2\kappa/D$ in total.
We can justify the second statement similarly. By the first part, one can cover $S_{1,D_{0}}$ by $(C_{0}D_{0}/\sqrt{m})^{m}$ balls of radius $2\kappa/D_{0}$. We then cover these balls by smaller balls of radius $2\kappa/D$, the number of such small balls is at most $(O(D/D_{0}))^{m}$. Thus there are at most $(O(D/\sqrt{m}))^{m}$ balls in total. ∎
4. Overcrowding estimate: proof of Theorem 2.6
We first introduce an important lemma (whose proof is delayed to the end of the section) to control the distance of a random vector with non i.i.d. entries to a subspace.
Lemma 4.1.
Let ${\mathbf{v}}\in{\mathbb{R}}^{n}$ be a random vector whose entries are independent (not necessarily identically distributed) and ${\mathbf{v}}_{i}$ have mean zero, ${\operatorname{Var}}({\mathbf{v}}_{i})\geq 1$ and $|{\mathbf{v}}_{i}|\leq T$ for some parameter $T$ with probability one. Then if $P_{H^{\perp}}$ is a deterministic orthogonal projection in ${\mathbb{R}}^{n}$ onto a subspace of dimension $k$, there exist constants $C,c,c^{\prime}$ depending on $p$ such that for any $0<t<1/2,$
$\sup_{{\mathbf{u}}\in{\mathbb{R}}^{n}}{\mathbb{P}}(\|P_{H^{\perp}}{\mathbf{v}}-{\mathbf{u}}\|_{2}\leq tT\sqrt{k}-c^{\prime}T)\leq C\exp(-ct^{2}k)$
Proof.
(of Theorem 2.6) We follow the proof strategy in [Ngu18]. Observe that as $L_{n}+F$ is symmetric, $\sigma_{n-k+1}(L_{n}+F)\leq\frac{ck}{\sqrt{n}}$ is equivalent to the event that there exists an $i$ such that
$\frac{-ck}{\sqrt{n}}\leq\lambda_{i}\leq\lambda_{i-k+1}\leq\frac{ck}{\sqrt{n}}.$
We let $I:=[-ck/\sqrt{n},ck/\sqrt{n}]$. Let us assume that $\lambda_{i}\in I$ and ${\mathbf{v}}_{i}$ is such that $(L_{n}+F){\mathbf{v}}_{i}=\lambda_{i}{\mathbf{v}}_{i}$. As $L_{n}+F$ is symmetric, and so the eigenvectors $v_{i}$ are orthogonal, then
$\left\|(L_{n}+F){\mathbf{v}}_{i}\right\|_{2}\leq\frac{ck}{\sqrt{n}}$ (11)
Write ${\mathbf{v}}_{j}=(v_{j1},\ldots,v_{jn})^{T}$ and let ${\mathbf{c}}_{i}$ denote the $i^{\text{th}}$ column of $L_{n}+F$. Let $V$ be the $k\times n$ matrix formed by the row vectors ${\mathbf{v}}_{i}^{T}$ and let ${\mathbf{w}}_{1},\dots,{\mathbf{w}}_{n}$ be its columns. Clearly, (11) is equivalent to
$\left\|\sum_{j=1}^{n}v_{ij}{\mathbf{c}}_{j}\right\|_{2}\leq\frac{ck}{\sqrt{n}}\text{\quad for all $1\leq i\leq k$}.$
For $J\subset[n]$, we use $V_{J}$ to indicate the matrix formed by the columns ${\mathbf{w}}_{j}$ with $j\in J$. Oberve that by construction,
$\sum_{1\leq j\leq n}\|{\mathbf{w}}_{j}\|_{2}^{2}=k.$
We now show that there exists a well-conditioned minor of submatrix of $V$. By Lemma 3.1, for any $11\leq l\leq k-$ 1, there exist distinct indices $i_{1},\dots,i_{l}$ such that
$\sigma_{l}(Z_{i_{1},\dots,i_{l}})\geq c\max_{r\in\{l+1,\cdots,k\}}\sqrt{\frac{(r-l)\sum_{i=r}^{n}\sigma_{i}(V)^{2}}{nr}}\geq c^{\prime}\sqrt{\frac{(k-l)^{2}}{nk}}$ (12)
for some constant $c^{\prime}$, where the final inequality follows from setting $r=\lceil(l+k)/2\rceil$ and noting that $\sigma_{j}(V)=1$ for $j\in[k]$.
For notational convenience, we let $Y$ be the $l\times k$ matrix $Z_{(i_{1},\ldots,i_{l})}^{T}$ and $X$ be a right inverse of $Y$ so that $YX=I_{l}$. We define
$A=({\mathbf{c}}_{i_{1}},\ldots,{\mathbf{c}}_{i_{l}})Y+({\mathbf{c}}_{i_{l+1}},\ldots,{\mathbf{c}}_{i_{n}})Y^{\prime},$ (13)
where $Y^{\prime}=Z_{(i_{l+1},\ldots,i_{n})}^{T}$. Multiplying (13) on the right by $X$ yields
$AX=({\mathbf{c}}_{i_{1}},\ldots,{\mathbf{c}}_{i_{l}})+({\mathbf{c}}_{i_{l+1}},\ldots,{\mathbf{c}}_{i_{n}})Y^{\prime}X.$
Recall that by assumption, each column of $A$ has norm bounded by $ck/\sqrt{n}$ and by (12),
$\|X\|\ll\sqrt{\frac{kn}{(k-l)^{2}}}.$
Thus,
$\|AX\|_{HS}\leq\|A\|_{HS}\|X\|\ll\frac{k^{2}}{(k-l)}.$ (14)
Let $H$ denote the subspace spanned by $({\mathbf{c}}_{i_{l+1}},\ldots,{\mathbf{c}}_{i_{n}})$. (14) implies that
${\operatorname{dist}}({\mathbf{c}}_{i_{1}},H)^{2}+\ldots+{\operatorname{dist}}({\mathbf{c}}_{i_{l}},H)^{2}\ll\frac{k^{4}}{(k-l)^{2}}.$
Thus, we have reduced the event in (11) to the even that
${\operatorname{dist}}({\mathbf{c}}_{i_{1}},H)=O\left(\frac{k^{2}}{(k-l)}\right)\wedge\cdots\wedge{\operatorname{dist}}({\mathbf{c}}_{i_{l}},H)=O\left(\frac{k^{2}}{(k-l)}\right)$
which we denote by $\mathcal{E}_{i_{1},\dots,i_{l}}$. However, due to the symmetry and structure of the Laplacian, these subevents are related in a complicated way. To begin decoupling the events, we make the following observation. For any $I\subset[n]$,
${\operatorname{dist}}({\mathbf{c}}_{i},H)\geq{\operatorname{dist}}({\mathbf{c}}_{i,I},H_{I}),$ (15)
where $c_{i,I}$ and $H_{I}$ are the projections of ${\mathbf{c}}_{i}$ and $H$ onto the coordinates indexed by $I$. Without loss of generality, we assume that $(i_{1},\ldots,i_{l})=(1,\ldots,l)$. We utilize (15) to note that $\mathcal{E}_{1,2,\dots,l}$ implies the event
$\mathcal{F}_{1,\dots,l}$ $:=\left({\operatorname{dist}}({\mathbf{c}}_{1,\{l,\dots,n\}},H_{k,\dots,n}=O\left(\frac{k^{2}}{(k-l)}\right)\right)\wedge\dots\wedge\left({\operatorname{dist}}({\mathbf{c}}_{l,\{l,\dots,n\}},H_{k,\dots,n}=O\left(\frac{k^{2}}{(k-l)}\right)\right)$
Note that now we have that ${\mathbf{c}}_{i,\{l,\dots,n\}}$ is independent of ${\mathbf{c}}_{j,\{l,\dots,n\}}$ for $i,j\leq l$ and $i\neq j$. However, due to the structure of the Laplacian, ${\mathbf{c}}_{i,\{l,\dots,n\}}$ is not independent of $H_{l,\dots,n}$
We now set $l=k/2$ and we introduce an integer parameter $\tau$, which will be a sufficiently large constant depending on $c$ and $p$, and divide our indices $1,\dots,l$ into $l/2\tau$ sets,
$J_{1}=\{1,\dots,2\tau\},J_{2}=\{2\tau+1,4\tau\},\dots,J_{l/2\tau}=\{l-2\tau+1,\dots,l\}$
As ${\mathbf{c}}_{i}$ is a column of $L_{n}+F$,
${\mathbf{c}}_{i,\{l,\dots,n\}}={\mathbf{d}}_{i,\{l,\dots,n\}}+{\mathbf{f}}_{1,\{l,\dots,n\}}.$
We define two vectors for every $i\in[l/\tau]$,
${\mathbf{c}}_{i}^{*}=\sum_{j=(i-1)\tau+1}^{i\tau/2}{\mathbf{c}}_{j}=\sum_{j=(i-1)\tau+1}^{i\tau/2}{\mathbf{d}}_{j}+{\mathbf{f}}_{j}:={\mathbf{d}}_{i}^{*}+{\mathbf{f}}_{i}^{*}$
and
${\mathbf{c}}_{i}^{**}=\sum_{j=i\tau/2+1}^{i\tau}{\mathbf{c}}_{j}={\mathbf{d}}_{i}^{**}+{\mathbf{f}}_{i}^{**}.$
Every coordinate of $\bar{{\mathbf{d}}}_{i,\{l,\dots,n\}}:={\mathbf{d}}_{i,\{l,\dots,n\}}^{*}+{\mathbf{d}}_{i,\{l,\dots,n\}}^{**}$ is an i.i.d. binomial random variable, For a random variable $X\sim B(2\tau,p)$, by Chernoff’s bound, for any $10<\delta<$ 1,
${\mathbb{P}}(|X-2\tau p|\geq\delta 2\tau p)\leq 2\exp(-2\delta^{2}\tau p/3).$
If we set $\delta=1/2$ and $\tau=K\log n/k$ for a sufficiently large constant $K$ then
${\mathbb{P}}(|X-2\tau p|<\delta 2\tau p)\geq(1-k/5n).$
Let $D_{i}:=\{j\in\{l,\dots,n\}:(\bar{{\mathbf{d}}}_{i,\{l,\dots,n\}})_{j}\in[(1-\delta)2\tau p,(1+\delta)2\tau p]\}$. As the entries are independent, by Chernoff’s bound again,
${\mathbb{P}}(|D_{i}|\leq(1-k/5n)(n-l))\leq\exp(-kp/10).$ (16)
Now for $i\in[l/2\tau]$, we introduce the following auxiliary randomness inspired again by the “switching” method. We define a new random variable $\hat{{\mathbf{c}}}^{*}_{i}$ where $(\hat{{\mathbf{c}}}_{i}^{*})_{j}=({\mathbf{c}}^{*}_{i})_{j}$ for $j\notin D_{i}$ and for $j\in D_{i}$,
$(\hat{{\mathbf{c}}}^{*}_{i})_{j}=\begin{cases}&(\hat{{\mathbf{c}}}_{i}^{*})_{j}\text{ with probability }1/2\\ &({\mathbf{d}}_{i}^{**})_{j}+{\mathbf{f}}_{i}^{*}\text{ with probability }1/2\end{cases}$
and similarly, we define a new random variable $\hat{{\mathbf{c}}}^{**}_{i}$ where $(\hat{{\mathbf{c}}}_{i}^{**})_{j}=({\mathbf{c}}^{**}_{i})_{j}$ for $j\notin D_{i}$ and for $j\in D_{i}$,
$(\hat{{\mathbf{c}}}^{**}_{i})_{j}=\begin{cases}&(\hat{{\mathbf{c}}}_{i}^{**})_{j}\text{ with probability }1/2\\ &({\mathbf{d}}_{i}^{*})_{j}+{\mathbf{f}}_{i}^{**}\text{ with probability }1/2.\end{cases}$
We observe that $\hat{{\mathbf{c}}}_{i,\{l,\dots,n\}}^{*},H_{l,\dots,n}$ has the same joint distribution as ${\mathbf{c}}_{i,\{l,\dots,n\}}^{*},H_{l,\dots,n}$. Furthermore, the key benefit of this construction is that even upon conditioning on the outcome of $H_{l,\dots,n}$, ${\mathbf{c}}_{i}$, and ${\mathbf{c}}_{i}^{*}$, $\hat{{\mathbf{c}}}_{i}^{*}$ is sufficiently random with independent entries for $j\in I_{i}$.
Next, we note that by the triangle inequality, $\mathcal{F}_{1,\dots,l}$ implies the event
$\mathcal{G}_{1,\dots,l}:=\left({\operatorname{dist}}(\hat{{\mathbf{c}}}^{*}_{1,\{l,\dots,n\}},H_{l,\dots,n})=O\left(\frac{\tau k^{2}}{(k-l)}\right)\right)\wedge\dots\wedge\left({\operatorname{dist}}(\hat{{\mathbf{c}}}^{*}_{l/\tau,\{l,\dots,n\}},H_{l,\dots,n})=O\left(\frac{\tau k^{2}}{(k-l)}\right)\right).$
[…]
${\mathbb{P}}(\mathcal{E}^{c})\leq(l/2\tau\alpha)\exp(-kp\alpha/10)\leq\exp(-kp\alpha/20).$ (17)
On the event $\mathcal{E}$, we have at least $l/2\tau\alpha$ good indices. For any good index $i$, by (15),
${\mathbb{P}}\left({\operatorname{dist}}(\hat{{\mathbf{c}}}^{*}_{i,D_{i}},H_{l,\dots,n})\leq O\left(\frac{\tau k^{2}}{(k-l)}\right)\right)$ $\leq{\mathbb{P}}\left({\operatorname{dist}}(\hat{{\mathbf{c}}}^{*}_{i,D_{i}},H_{D_{i}})=O\left(\frac{\tau k^{2}}{(k-l)}\right)\right)$
$\leq C\exp(-\Theta(k))$
where the last inequality follows from Lemma 4.1 and $k$ sufficiently large. Note that our choice of $l$ and size of $|D_{i}|$ guarantees that the corank of $H_{D_{i}}$ is at least $k/3$. Therefore, by independence of the randomness on the good indices, we have that
${\mathbb{P}}(\mathcal{E}_{1,\dots,l})$ $\leq{\mathbb{P}}(\mathcal{E}_{1,\dots,l}|\mathcal{E})+{\mathbb{P}}(\mathcal{E}^{c})$
$\leq(C\exp(-\Theta(k)))^{l/2\tau\alpha}+\exp(-kp\alpha/20)\leq C\exp(-\Theta(k^{2}/\alpha\log(n/k)))+\exp(-k\alpha/20).$
Finally, we set $\alpha=\sqrt{k/\log n}$. ∎
It remains to prove the distance result.
Proof.
(of Lemma 4.1) We show that $f({\mathbf{x}})=\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}$ is convex and 1-Lipschitz. For convexity, we observe that for $0\leq s\leq 1$,
$\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}$ $=\|P_{H^{\perp}}s{\mathbf{x}}-s{\mathbf{u}}+P_{H^{\perp}}(1-s){\mathbf{x}}-(1-s){\mathbf{u}}\|_{2}$
$\leq s\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|+(1-s)\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}.$
Next, note that ${\mathbf{x}},{\mathbf{y}}\in{\mathbb{R}}^{n}$,
$\Big{|}\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}-\|P_{H^{\perp}}{\mathbf{y}}-{\mathbf{u}}\|_{2}\Big{|}$ $\leq\|P_{H^{\perp}}({\mathbf{x}}-{\mathbf{y}})\|_{2}$
$\leq\|{\mathbf{x}}-{\mathbf{y}}\|_{2}$
so that $f({\mathbf{x}})$ is 1-Lipschitz. Thus, by Talagrand’s inequality, for absolute constants $K,\kappa>0$ and any $\lambda>0$,
${\mathbb{P}}(|f({\mathbf{x}})-M(f({\mathbf{x}})|\geq\lambda T)\leq K\exp(-\kappa\lambda^{2})$ (18)
where $M(f({\mathbf{x}}))$ is a median of $f({\mathbf{x}})$. We use a second moment argument to control the median. We first calculate that
${\mathbf{E}}\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}^{2}={\mathbf{E}}{\mathbf{x}}^{T}P_{H^{\perp}}{\mathbf{x}}-2{\mathbf{E}}{\mathbf{x}}^{T}P_{H^{\perp}}{\mathbf{u}}+\|{\mathbf{u}}\|^{2}=\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}+\|{\mathbf{u}}\|_{2}^{2}$
where $p_{ij}:=(P_{H^{\perp}})_{ij}$.
$Y$ $=\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}^{2}-{\mathbf{E}}\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}^{2}$
$={\mathbf{x}}^{T}P_{H^{\perp}}{\mathbf{x}}-2{\mathbf{x}}^{T}P_{H^{\perp}}{\mathbf{u}}+\|{\mathbf{u}}\|^{2}-\sum_{i=1}^{n}p_{ii}{\mathbf{E}}x_{i}^{2}-\|{\mathbf{u}}\|_{2}^{2}$
$=\sum_{ij}p_{ij}(x_{i}x_{j}-\delta_{ij})-2\sum_{i=1}^{n}x_{i}(P_{H^{\perp}}{\mathbf{u}})_{i}.$
Therefore,
${\mathbf{E}}Y^{2}$ $={\mathbf{E}}\left(\sum_{ij}p_{ij}(x_{i}x_{j}-\delta_{ij})\right)^{2}-4{\mathbf{E}}\left(\sum_{i=1}^{n}x_{i}(P_{H^{\perp}}{\mathbf{u}})_{i}\right)\left(\sum_{ij}p_{ij}(x_{i}x_{j}-\delta_{ij})\right)+4{\mathbf{E}}\left(\sum_{i=1}^{n}x_{i}(P_{H^{\perp}}{\mathbf{u}})_{i}\right)^{2}$
$\leq O\left(T^{4}\sum_{ij}p_{ij}^{2}+T^{4}\sqrt{\left(\sum_{i}(P_{H^{\perp}}{\mathbf{u}})_{i}^{2}\right)\left(\sum_{ij}p_{ij}^{2}\right)}\right)$
$=\left(T^{4}k+T^{2}\|{\mathbf{u}}\|_{2}^{2}k\right)$
where we have invoked the fact that $\sum_{ij}p_{ij}^{2}=\sum_{i}p_{ii}=k$.
Since we make no assumptions on $\|{\mathbf{u}}\|_{2}$, we need to divide into two cases.
Case 1. We assume that $(1-2t)T\sqrt{\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}}\leq\|{\mathbf{u}}\|_{2}\leq(1+2t)T\sqrt{\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}}$, then we have that
${\mathbf{E}}Y^{2}=O(T^{4}k).$
By Markov’s inequality, the median of $|Y|$ is $O(T^{2}\sqrt{k})$, which implies that the median of $\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}^{2}$ is at least $\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}+\|u\|_{2}^{2}-O(T^{2}\sqrt{k})\geq T^{2}k+\|{\mathbf{u}}\|_{2}^{2}-O(T^{2}\sqrt{k})$. We can therefore deduce that the median of $\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}$ is at least $\sqrt{T^{2}k+\|{\mathbf{u}}\|_{2}^{2}-O(T^{2}\sqrt{k})}=\sqrt{T^{2}k+\|{\mathbf{u}}\|_{2}^{2}}-O(T^{2})$. Returning to Talagrand’s inequality, (18), we find that
${\mathbb{P}}(\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}\leq tT\sqrt{k}-O(T^{2}))$ $\leq{\mathbb{P}}(\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}\leq\sqrt{T^{2}k+\|{\mathbf{u}}\|_{2}^{2}}tT\sqrt{k}-O(T^{2})-Tt\sqrt{k})$
$\leq{\mathbb{P}}(|\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}-M(\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2})|\geq Tt\sqrt{k})$
$\leq C\exp(-c^{\prime}t^{2}k).$
Case 2. Now we consider the case where $\|{\mathbf{u}}\|_{2}\leq(1-2t)T\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$ or $\|{\mathbf{u}}\|_{2}\geq(1+2t)T\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$. Then, if $\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}\leq t\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$ then by the triangle inequality, either $\|P_{H^{\perp}}{\mathbf{x}}\|_{2}\leq(1-t)T\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$ or $\|P_{H^{\perp}}{\mathbf{x}}\|_{2}\geq(1+t)T\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$ which reduces to Case 1. ∎
5. No-gaps delocalization: proof of Theorem 2.3 for the approximated vectors
First, recall the Erdős-Littlewood-Offord result from [NW24].
Theorem 5.1.
Let ${\mathbf{w}}=(w_{1},\dots,w_{N})\in{\mathbb{R}}^{N}$.
  • Assume that at least $N^{\prime}$ elements of the $w_{i}$ satisfy $|w_{i}|\geq r$. Let $X=(x_{1},\dots,x_{N})$ where $x_{i}$ are iid copies of a random variable $\xi$ of mean zero, variance one, and bounded $(2+\varepsilon)$-moment. Then
    $\rho_{r}({\mathbf{w}}):=\sup_{a\in{\mathbb{R}}}{\mathbb{P}}(|X\cdot{\mathbf{w}}-a|<r)=O(\frac{1}{\sqrt{N^{\prime}}}).$
    Consequently, if
    for every $\text{for every }I\subset[N]\text{ with }|I|\geq N-N^{\prime}\text{ we have }\|{\mathbf{w}}_{I}\|_{2}^{2}\geq|I|r^{2},$ (19)
    then the above holds.
  • Furthermore, if we assume that for all $w\in{\mathbb{R}}$, the vector $\operatorname{supp}({\mathbf{w}}-w\mathbf{1})$ has at least $N^{\prime}$ components that have absolute values at least $r$. Then for $X=(x_{1},\dots,x_{N})$ with $x_{i},$ for $1\leq i\leq N$, being i.i.d. copies of $\xi$ we have the following affine analog
    $\rho_{L,r}({\mathbf{w}}):=\sup_{a,w\in{\mathbb{R}}}{\mathbb{P}}(|X\cdot({\mathbf{w}}-w\mathbf{1})-a|<r)=O(\frac{1}{\sqrt{N^{\prime}}}).$
    As a consequence, for any $R>r$
    $\rho_{L,R}({\mathbf{w}})=\sup_{a,w\in{\mathbb{R}}}{\mathbb{P}}(|X\cdot({\mathbf{w}}-w\mathbf{1})-a|<R)=O(\frac{R}{r\sqrt{N^{\prime}}}).$ (20)
    Also, if for any $w$
    $\|({\mathbf{w}}-w\mathbf{1})_{I}\|_{2}^{2}\geq|I|r^{2}\text{ for any }I\text{ such that }|I|\geq N-N^{\prime},$ (21)
    then the above conclusion holds.
Here the implied constants depend on $\xi$.
Note that for Theorem 2.3 it suffices to assume
$\frac{1}{\log\log n}\leq\lambda_{n}<1/2.$
It what follows we will mainly work with $\overline{L}_{n}$, the proof for $\overline{L}^{\prime}_{n}$ (or for $\overline{L}^{\prime}_{n-1}$) is identical.
5.1. Decomposition of vectors
Let
$\delta_{n}:=(\log n)^{-\frac{A_{0}}{\lambda_{n}}},\mbox{ where $A_{0}$ to be chosen sufficiently large.}$. (22)
First, we let $I$ be the set of first $\lambda_{n}n$ indices (in the end we will take union bound over $I$) and will bound the probability of the following event $\mathcal{E}$: there exists a ${\mathbf{v}}\in(\frac{1}{D}{\mathbb{Z}})^{n}$ satisfying (7) whose first $\lceil\lambda_{n}n\rceil$ coordinates are the same (in the $l_{2}$-norm), that is for some $a$
$\|{\mathbf{v}}_{I}-a\mathbf{1}/\sqrt{n}\|_{2}\leq\delta_{n}$
and
$\|(\overline{L}_{n}-\lambda_{0}){\mathbf{v}}\|_{2}=O(\sqrt{n}\log n\times n^{-C}).$
Let
$k:=k_{n}=\lfloor\lambda_{n}n/4\rfloor.$
We partition the index set $[\lceil\lambda_{n}n\rceil+1,n]$ into subsequences $J_{1},\dots,J_{\ell}$ of consecutive numbers so that $|J_{i}|=k$ for $i<\ell$ and $k\leq|J_{\ell}|<2k$.
We have
$k\geq\lambda_{n}n/8\quad\textrm{and}\quad\ell<8\lambda_{n}^{-1}.$
We let $t_{0}>1$ be a jumping parameter to be chosen later, for instance we can choose
$t_{0}=\log^{64}n.$ (23)
Also, let
$\alpha_{n}=\log^{-2}n.$ (24)
Given ${\mathbf{v}}$, we will partition $[n]$ into two subsets, $I_{m}$ (mixed) and $I_{s}$ (sparse), where $I_{s}$ is the union of the intervals $J_{i_{0}},J_{{i_{1}}},\dots$ above which can be defined step by step below: at each step $i\geq 0$, find a ${\mathbf{v}}_{J_{i}}$ such that there exists $a_{i}\in{\mathbb{Z}}/D$ and a subset $J_{i}^{\prime}\subset J_{i}$ of size at least $|J_{i}|-\alpha_{n}\lambda_{n}n$ which satisfies
$\|{\mathbf{v}}_{J_{i}^{\prime}}-(a_{i}\mathbf{1}/\sqrt{n})_{J_{i}^{\prime}}\|_{2}\leq\delta_{n}t_{0}^{i}.$ (25)
Hence $I_{s}$ is associated with a maximum index $0\leq i_{0}\leq 8\lambda_{n}^{-1}$ and a partial ordering of $\{1,\dots,\lfloor\lambda_{n}^{-1}\rfloor\}$ where the $J_{i}$ appeared in each step $i$.
We first pause for a few remarks.
Remark 5.2.
Each ${\mathbf{v}}$ will also gives rise to a parameter $0\leq i_{0}\leq 8\lambda_{n}^{-1}$ such that all of the intervals from $I_{s}$ are $\delta_{n}t_{0}^{{i_{0}}}$-closed to some $a_{i}\mathbf{1}/\sqrt{n}$, while in $I_{m}$, if $J_{i}$ does not belong to $I_{s}$ then by definition, for any $a\in{\mathbb{Z}}/D$ and any $J_{i}^{\prime}\subset J_{i}$ such that $|J_{i}^{\prime}|\geq|J_{i}|-\alpha_{n}\lambda_{n}n$ we have
$\|{\mathbf{v}}_{J_{i}^{\prime}}-(a_{i}\mathbf{1}/\sqrt{n})_{J_{i}^{\prime}}\|_{2}>\delta_{n}t_{0}^{i_{0}+1}.$
This implies that for any $a$, there are at least $\alpha_{n}\lambda_{n}n$ entries $v_{i}\in{\mathbf{v}}_{J_{i}}$ such that
$|v_{i}-a/\sqrt{n}|\geq\frac{\delta_{n}t_{0}^{i_{0}+1}}{\sqrt{|J_{i}|}}=\frac{\delta_{n}t_{0}^{i_{0}+1}\lambda_{n}^{-1/2}}{\sqrt{n}}.$
In particularly, Condition 21 of Theorem 5.1 is satisfied, and so
$\rho_{L,\frac{\delta_{n}t_{0}^{i_{0}+1}\lambda_{n}^{-1/2}}{\sqrt{n}}}({\mathbf{v}}_{J_{i}})=O(\frac{1}{\sqrt{\alpha_{n}\lambda_{n}n}}).$ (26)
For convenience set
$r_{n}:=\frac{\delta_{n}t_{0}^{i_{0}+1/2}\lambda_{n}^{-1/2}}{\sqrt{n}},$ (27)
with a saving factor $t_{0}^{1/2}$ for later use.
Remark 5.3.
Note that for each $i\leq i_{0}$,
$t_{0}^{i}\delta_{n}\leq t_{0}^{i_{0}}\delta_{n}<r_{n}\sqrt{|J_{i}|}.$
So for each fixed $a$, the set $\{{\mathbf{v}}_{J_{i}}\in{\mathbb{R}}^{|J_{i}|},\|{\mathbf{v}}_{J_{i}}-a\mathbf{1}/\sqrt{n}\|_{2}\leq t_{0}^{i}\delta_{n}\}$ trivially accepts an $r_{n}\sqrt{J_{i}}$-net of size 1.
As a consequence, the set $\{{\mathbf{v}}_{J_{i}}\in{\mathbb{R}}^{|J_{i}|},\|{\mathbf{v}}_{J_{i}}-a\mathbf{1}/\sqrt{n}\|_{2}\leq t_{0}^{i}\delta_{n},\mbox{ for some }a\in{\mathbb{Z}}/D,|a|\leq\sqrt{n}\}$, trivially accepts an $r_{n}\sqrt{J_{i}}$-net of size $O(D\sqrt{n})$.
We let $I_{m}=[n]\setminus I_{s}$. We write ${\mathbf{v}}_{m}:={\mathbf{v}}_{I_{m}}$ and ${\mathbf{v}}_{s}:={\mathbf{v}}_{I_{s}}$. Note that ${\mathbf{v}}_{s}$ can be $r_{n}\sqrt{|I_{s}|}$-approximated by a vector ${\mathbf{v}}_{s}^{\prime}$ which is a concatenation of vectors of form $(a_{i}\mathbf{1}/\sqrt{n})_{J_{i}^{\prime}}$. We let ${\mathbf{v}}^{\prime}=({\mathbf{v}}_{s}^{\prime},{\mathbf{v}}_{m}^{\prime})$ where ${\mathbf{v}}_{m}^{\prime}={\mathbf{v}}_{m}$, and also, basing on Subsection 2.4, without loss of generality we can assume the entries of ${\mathbf{v}}_{m}^{\prime}$ (which is the subvector ${\mathbf{v}}_{J_{i}}$ where there is no $a_{i}$, and from the subvectors ${\mathbf{v}}_{J_{i}\backslash J_{i}^{\prime}}$ where there exists $a_{i}$) are from ${\mathbb{Z}}/D$.
Note that by defition,
$\|(\overline{L}_{n}-\lambda_{0}){\mathbf{v}}^{\prime}\|_{2}=O(\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|});\mbox{ and }|\sum_{i}v_{i}^{\prime}|\leq r_{n}|I_{s}|;\mbox{ and }\|{\mathbf{v}}^{\prime}\|_{2}=1+O(r_{n}\sqrt{I_{s}}),$ (28)
where $\sqrt{n}\log n$ comes from the upper bound for $\|L-\lambda_{0}\|_{2}$, and by triangle inequality
$\|(\overline{L}_{n}-\lambda_{0}){\mathbf{v}}^{\prime}\|_{2}\leq\|(L-\lambda_{0}){\mathbf{v}}\|_{2}+\|(L-\lambda_{0})({\mathbf{v}}-{\mathbf{v}}^{\prime})\|_{2}=O(\sqrt{n}\log n\times n^{-C})+O(\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|}).$ (29)
Although we will mention in more details later, as of this point we remark that our main starting point is to pass the event considered in Theorem 2.3 to the event that there exists ${\mathbf{v}}^{\prime}$ such that (29) holds.
As of this point, roughly speaking we have extended the special vector ${\mathbf{v}}_{I}$ to a longest possible vector ${\mathbf{v}}_{s}$ whose entries take only a few values, while ${\mathbf{v}}_{m}$ is the left-over that we no longer detect entries with high multiplicities.
Case 1. Let $\mathcal{E}_{1}$ be the event that there exists a non-zero ${\mathbf{v}}\in(\frac{1}{D}{\mathbb{Z}})^{n}$ satisfying (7) whose first $\lceil\lambda_{n}n\rceil$ coordinates are the same, and the resulting $I_{m}$ (from ${\mathbf{v}}$) is empty in the sense that for all $i$ we can approximate ${\mathbf{v}}_{J_{i}}$. It is crucial to note that in this case we don’t have to take union bound over $I$ (of Theorem 2.3) because the event $I_{m}$ being empty does not depend on $I$.
We let ${\mathbf{v}}^{\prime}$ be obtained from ${\mathbf{v}}$ by replacing ${\mathbf{v}}_{J_{i}^{\prime}}$ with $a_{i}\mathbf{1}/\sqrt{n}$, and together with some $v_{j}\in{\mathbb{Z}}/D$ for each $j\in J_{i}\backslash J_{i}^{\prime}$. As every sub-interval belongs to a level set, note that in this case we have
$\|{\mathbf{v}}_{J_{i}^{\prime}}-a_{i}\mathbf{1}/\sqrt{n}\|_{2}\leq\delta_{n}t_{0}^{i_{0}}\leq\delta_{n}t_{0}^{8\lambda_{n}^{-1}}.$
The factor $t_{0}^{8\lambda_{n}^{-1}}$ can be as large as $(\log n)^{512\lambda_{n}^{-1}}$, but as $\delta_{n}$ is sufficiently small from (22), the RHS is still small.
We will consider a $\delta_{n}t_{0}^{\lambda_{n}^{-1}}$-approximation of the vectors. There are at most $D^{\alpha_{n}n}\times D^{\lambda_{n}^{-1}}$ choices of ${\mathbf{v}}^{\prime}$ that can arise with empty $I_{m}$ with the first factor stands for the amount of choices of the outliers and the second factors to be the choices of the representation of each interval (assuming that $1/D$ is small compared to $\delta_{n}t_{0}^{\lambda_{n}^{-1}}$, which is necessary for the approximation to make sense).
We write ${\mathbf{v}}^{\prime}=({\mathbf{v}}_{1}^{\prime},{\mathbf{v}}_{2}^{\prime})$ where ${\mathbf{v}}_{1}^{\prime}$ are from the $(a_{i}/\sqrt{n})$ and the support of ${\mathbf{v}}_{2}^{\prime}$ is at most $\alpha_{n}n$. It is important to emphasize that we are working with the event that there exists ${\mathbf{v}}^{\prime}$ of that form such that $(L-\lambda_{0}){\mathbf{v}}^{\prime}$ is small.
Subcase 1.1. Assume first that the support of the $\{a_{i},1\leq i\leq\ell\}$ has length at least $2\delta_{n}^{\prime}$, where
$\delta_{n}^{\prime}:=8\lambda_{n}^{-3/2}(\delta_{n}t_{0}^{\lambda_{n}^{-1}})\sqrt{\log n}.$ (30)
Then either the smallest or largest values of $a_{i}$ is $\delta_{n}^{\prime}$-far from $\lambda_{n}^{-1}/2$ of the rest. Assume without loss of generality that $a$, the multiplicity of the first segment, is $\delta_{n}^{\prime}$-far from $\lambda_{n}^{-1}/2$ of the multiplicities $j_{0}$, since the level set comes together with a partial order and $j_{0}$ covers every others. Consider the event
$((\overline{L}_{n}-\lambda_{0}){\mathbf{v}}^{\prime})_{i}\leq 2\sqrt{\log n}(\delta_{n}t_{0}^{8\lambda_{n}^{-1}}),\mbox{where $i\in I_{j_{0}}$}.$
This can be written as (here $x_{ij}$ are the entries of the adjacency matrix)
$\sum_{j\neq i}\frac{a_{j}}{\sqrt{n}}x_{ij}-\frac{a_{i}}{\sqrt{n}}(\sum_{j\neq i}x_{ij})-(\lambda_{0}\sum_{i}v_{i}^{\prime})-f$ $=x_{i1}(a-a_{i})/\sqrt{n}+x_{i2}(a-a_{i})/\sqrt{n}+\dots$
$=(x_{i1}+x_{i2}+\dots+x_{i|J_{1}|})(a-a_{j_{0}})/\sqrt{n}+\dots,$
where $f$ depends on the $a_{i}$ but not on the $x_{ij}$  [^5] This $f$ comes from the part ${\mathbf{E}}L_{n}$ in the definition of $\overline{L}_{n}$..
Note that there is also the part for ${\mathbf{v}}_{2}^{\prime}$ but we can fix the randomness over its vector support.
Hence for all $i\in I_{j_{0}}$, because the sum $x_{i1}+x_{i2}+\dots+x_{i|J_{1}|}$ takes value in [±$[\pm\Theta(\sqrt{\lambda_{n}n})]$ by Central Limit Theorem,
${\mathbb{P}}(((\overline{L}_{n}-\lambda_{0}){\mathbf{v}}^{\prime})_{i}\leq 2\sqrt{\log n}(\delta_{n}t_{0}^{8\lambda_{n}^{-1}})\leq 1/2,$
provided that $\delta_{n}^{\prime}\sqrt{\lambda_{n}}$ is sufficiently larger than $\sqrt{\log n}(\delta_{n}t_{0}^{\lambda_{n}^{-1}})$, which is satisfied by (30).
Taking over all $i\in\cup_{i}I_{j_{0}}$, by independence we hence obtain
${\mathbb{P}}\Big{(}((\overline{L}_{n}-\lambda_{0}){\mathbf{v}}^{\prime})_{j}\leq 2\sqrt{\log n}(\delta_{n}t_{0}^{8\lambda_{n}^{-1}}),i\in\cup_{i}I_{j_{0}}\Big{)}\leq(1/2)^{(n/2-\alpha_{n}n)}.$
Therefore, by tensorization (Lemma 3.3), we hence obtain
${\mathbb{P}}\Big{(}\|(\overline{L}_{n}-\lambda_{0}){\mathbf{v}}^{\prime}\|_{2}=O(\sqrt{\log n}(\delta_{n}t_{0}^{8\lambda_{n}^{-1}})\sqrt{n})\Big{)}\leq(1/2)^{n/3}.$
Taking union bound over $D^{\alpha_{n}n}\times D^{\lambda_{n}^{-1}}\leq n^{Cn/(\log n)^{2}}$ choices of ${\mathbf{v}}^{\prime}$, we see that
${\mathbb{P}}({\mathcal{E}}_{1})\leq(1/2)^{n/4}.$
Subcase 1.2. We next consider the case that all the values $a_{i}$ are of distance at most $2\delta_{n}^{\prime}$. Then there exists $a$ such that
$\|{\mathbf{v}}_{1}^{\prime}-a\mathbf{1}/\sqrt{n}\|_{2}\leq O(\delta_{n}^{\prime}).$
It thus follows that, as ${\mathbf{v}}$ is almost orthogonal to $\mathbf{1}$ because ${\mathbf{v}}$ is chosen as a non-trivial eigenvector (that is $\|{\mathbf{v}}-a\mathbf{1}/\sqrt{n}\|$ has order 1), $\|{\mathbf{v}}_{2}^{\prime}-a\mathbf{1}/\sqrt{n}\|_{2}$ is of order 1. By triangle inequality,
$\|(\overline{L}_{n}-\lambda_{0})({\mathbf{v}}^{\prime}-a\mathbf{1}/\sqrt{n}\|_{2}=O(\sqrt{\log n}(\delta_{n}t_{0}^{\lambda_{n}^{-1}})\sqrt{n})+\sqrt{n}\log n(\delta_{n}^{\prime})\ll\sqrt{n}/\log n,$
provided that $\delta_{n}$ is chosen as in (22).
On the other hand, by the Levy anti-concentration bound (see for instance [RV08, Lemma 2.6]), there exists $p<1$ such that
$\rho_{1/2}(\sum_{j\in\operatorname{supp}({\mathbf{v}}_{2}^{\prime})}x_{ij}(v_{j}^{\prime}-a/\sqrt{n}))<p.$
In particularly, by tensorization,
${\mathbb{P}}\Big{(}(\overline{L}_{n}-\lambda_{0})({\mathbf{v}}^{\prime}-a\mathbf{1}/\sqrt{n})=o(\sqrt{n})\Big{)}\leq p^{n-2\alpha_{n}n}.$
taking union bound over the choices of ${\mathbf{v}}^{\prime}$, we obtain $D^{\lambda_{n}^{-1}}D^{\alpha_{n}n}p^{n-2\alpha_{n}n}$, which is small as in Subcase 1.1.
We now move to the harder scenario.
Case 2. Let $\mathcal{E}_{2}$ be the event that there exists a non-zero ${\mathbf{v}}\in(\frac{1}{D}{\mathbb{Z}})^{n}$, almost orthogonal to $\mathbf{1}$, whose first $\lceil\lambda_{n}n\rceil$ coordinates are the same, $\|(\overline{L}_{n}-\lambda_{0}){\mathbf{v}}\|_{2}\leq n^{-C}$, and the resulting $I_{m}$ (from ${\mathbf{v}}$) is not empty. We have
$|I_{m}|\geq\lambda_{n}n/8\mbox{ and }|I_{s}|\geq\lambda_{n}n.$
Our treatment here is guided by the discussion from Subsection 2.5. We first pass to the approximated vector ${\mathbf{v}}^{\prime}$ and consider the event (29). Assume that $I_{s}$ consists of $i_{0}\leq\lambda_{n}^{-1}$ segments. We now describe a function $f$ from subsets of $[n]$ (that can occur as $I_{s}$) to subsets of $[n]$. We describe $F=f(I_{s})$, and write $I_{m}$ for $[n]\setminus I_{s}$, but note that $F$ does not depend on ${\mathbf{v}}$.
  • If $|I_{s}|>|I_{m}|$, we let $F$ be the first $|I_{m}|$ elements of $I_{s}$.
  • Otherwise, we let $J_{*}$ be the first $J_{i}$ in $I_{m}$ and we let $F$ be the union of $I_{s}$ and the first $|I_{m}|-|I_{s}|$ elements of $I_{m}$ that are not in $J_{*}$. (This is possible because $|J_{*}|\leq\lambda_{n}n/2\leq|I_{s}|.$)
The following diagram explains the construction with []delimited-[][-][ - ] denoting $I_{m},I_{s}$ and ()(-)( - ) denoting $F$:
[($F$)$I_{s}$][$I_{m}$]
Figure 2. $|I_{s}|>|I_{m}|$
[($I_{s}$][$F$)$I_{m}$]
Figure 3. $|I_{s}|<|I_{m}|$
In either case, we have $|F|=|I_{m}|$ and if $I_{m}$ is non-empty, it contains some $J_{i}$ that does not intersect $F$. Therefore, we create a square submatrix as well as an i.i.d. thin submatrix.
Let $c^{\ast}_{n}$ be chosen as
$c^{\ast}_{n}=\log^{-2}n$ (31)
Let ${\mathcal{E}}_{drop}$ be the event that there is a square submatrix $M_{A\times B}$ of $M=L_{n}$ of dimension $|A|=|B|$ at least $\lfloor n^{1-\delta_{0}}\rfloor$ for a sufficiently small constant $\delta_{0}$ so that $n^{1-\delta_{0}}<\lambda_{n}n$. with the $c^{\ast}_{n}|A|$-least singular value at least $c_{0}c^{\ast}_{n}\sqrt{|A|}$. Theorem 2.6 tells us that ${\mathbb{P}}({\mathcal{E}}_{drop})\leq e^{O(n)}e^{-\Theta(\lfloor n^{1-\delta_{0}}\rfloor)\sqrt{n}}$ (more specifically $e^{c^{\ast}_{n}n\log n}e^{-\Theta(c^{\ast}_{n}\lfloor n^{1-\delta_{0}}\rfloor)\sqrt{n}}$) and thus ${\mathbb{P}}({\mathcal{E}}_{drop})\leq e^{-c_{drop}n^{3/2-2\delta_{0}}},$ for some $c_{drop}>0$.
We wish to bound the probability of $\mathcal{E}_{2}\setminus{\mathcal{E}}_{drop}$. Since for ${\mathbf{v}}$ causing $\mathcal{E}_{2}$, we have $|I_{m}|>\lambda_{n}n$.
Outside of ${\mathcal{E}}_{drop}$, the square matrix $M_{F\times I_{m}}$ is near isometry: it has a rectangular submatrix $M_{12}=M_{F\times{{I_{m}^{\prime}}}}$ of dimension $|I_{m}|\times(|I_{m}|-c^{\ast}_{n}|I_{m}|)$ with the least (non-trivial) singular value at least $c_{0}c^{\ast}_{n}\sqrt{|I_{m}|}$.
Given subsets $S$ (for $I_{m}$), and $I_{m}^{\prime}\subset I_{m}$ of size $|I_{m}^{\prime}|=|I_{m}|-c^{\ast}_{n}|I_{m}|$, let $\mathcal{E}_{2.1}(S,{{{{I^{\prime}_{m}}}}}))$ be the event there exists ${\mathbf{v}}^{\prime}$ such that the above holds, and $\|(\overline{L}_{n}-\lambda_{0}){\mathbf{v}}^{\prime}\|_{2}$ is $O(\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|})$.
Lemma 5.4.
For all $S,{{{{I^{\prime}_{m}}}}}$ such that $\mathcal{E}_{2.1}(S,{{{{I^{\prime}_{m}}}}})$ is defined, we have
${\mathbb{P}}(\mathcal{E}_{2.1}(S,{{{{I^{\prime}_{m}}}}}))\leq\left(\frac{1}{t_{0}}\right)^{\lambda_{n}n/16}.$
Proof.
(of Lemma 5.4) We fix a choice of ${\mathbf{v}}_{s}^{\prime}$ and ${\mathbf{v}}_{I_{m}\setminus{{I^{\prime}_{m}}}}^{\prime}$. In what follows
$M_{11}=L_{F\times(I_{s}\cup(I_{m}\backslash I_{m}^{\prime}))},M_{12}=L_{F\times I_{m}^{\prime}},M_{21}=L_{[n]\backslash F\times(I_{s}\cup(I_{m}\backslash I_{m}^{\prime}))},M_{22}=L_{[n]\backslash F\times I_{m}^{\prime}}.$
[…]
Recall that the matrix $M_{12}$ is near isometry. Let $H_{12}$ be an $(|I_{m}|-c^{\ast}_{n}|I_{m}|)\times|I_{m}|$ matrix where $H_{12}M_{12}|_{I_{m}^{\prime}}=I_{I_{m}^{\prime}}$ (i.e. left inverse). By above (i.e. via. the application of Theorem 2.6), we have
$\|H_{12}\|_{2}\leq c_{0}^{-1}(c^{\ast}_{n})^{-1}/\sqrt{|I_{m}|}.$
Now we can rewrite $\|\overline{L}_{n}{\mathbf{v}}^{\prime}\|_{2}\leq\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|}$ as $\|M_{1}^{\prime}{\mathbf{v}}_{s}^{\prime\prime}+M_{2}^{\prime}{\mathbf{v}}_{m}^{\prime\prime}\|_{2}\leq\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|}$, where
${\mathbf{v}}_{s}^{\prime\prime}=({\mathbf{v}}_{s}^{\prime},{\mathbf{v}}_{I_{m}\backslash I_{m}^{\prime}}^{\prime});{\mathbf{v}}_{m}^{\prime\prime}={\mathbf{v}}_{I_{m}^{\prime}}^{\prime}$
as
$\|(M_{11}{\mathbf{v}}_{s}^{\prime\prime},M_{12}{\mathbf{v}}_{s}^{\prime\prime})+(M_{21}{\mathbf{v}}_{m}^{\prime\prime},M_{22}{\mathbf{v}}_{m}^{\prime\prime})\|_{2}\leq\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|}.$ (32)
In particularly, $\|M_{11}{\mathbf{v}}_{s}^{\prime\prime}+M_{12}{\mathbf{v}}_{m}^{\prime\prime}\|_{2}\leq\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|}$, and after applying $H_{12}$ we have that
$\|H_{12}M_{11}{\mathbf{v}}_{s}^{\prime\prime}+{\mathbf{v}}_{m}^{\prime\prime}\|_{2}\leq\sqrt{n}\log n\times r_{n}c_{0}^{-1}(c^{\ast}_{n})^{-1}\sqrt{|I_{s}|}/\sqrt{|I_{m}|}\leq\sqrt{n}\log n\times\lambda_{n}^{-1/2}r_{n}c_{0}^{-1}(c^{\ast}_{n})^{-1}(\frac{1-c_{m}}{c_{m}})^{1/2}$
$\leq\sqrt{n}\log n\times\lambda_{n}^{-1}r_{n}c_{0}^{-1}(c^{\ast}_{n})^{-1},$
where $|I_{m}|=c_{m}n$, with $\lambda_{n}/8<c_{m}<1-\lambda_{n}$. Here we note that $\delta_{n}$ is sufficiently small given all other parameters, and so the RHS bound will be small.
In other words, ${\mathbf{v}}_{m}^{\prime\prime}$ is almost determined if we fix $M_{11},M_{12}$ and ${\mathbf{v}}_{s}^{\prime\prime}$. As of this point, we will replace ${\mathbf{v}}_{m}^{\prime\prime}$ by ${\mathbf{u}}_{m}^{\prime\prime}=-H_{12}M_{11}{\mathbf{v}}_{s}^{\prime\prime}$. The latter vector is fixed if we condition on $M_{11},M_{12}$ and ${\mathbf{v}}_{s}^{\prime\prime}$. Note that
$\|{\mathbf{v}}_{m}^{\prime\prime}-{\mathbf{u}}_{m}^{\prime\prime}\|_{2}\leq\sqrt{n}\log n\times\lambda_{n}^{-1}r_{n}c_{0}^{-1}(c^{\ast}_{n})^{-1}=K_{n}r_{n},$
where
$K_{n}:=\sqrt{n}\log n\times\lambda_{n}^{-1}c_{0}^{-1}(c^{\ast}_{n})^{-1}.$
This approximation of $K_{n}r_{n}$ is rather big compared to $\sqrt{n}r_{n}$, but it is still smaller than $t_{0}^{1/2}\sqrt{n}r_{n}$ because $t_{0}$ is a much larger power of $\log n$.
We have
$\|M_{21}{\mathbf{v}}_{s}^{\prime\prime}+M_{22}{\mathbf{u}}_{m}^{\prime\prime}\|_{2}\leq\|M_{21}{\mathbf{v}}_{s}^{\prime\prime}+M_{22}{\mathbf{v}}_{m}^{\prime\prime}\|_{2}+\|M_{22}({\mathbf{v}}_{m}^{\prime\prime}-{\mathbf{u}}_{m}^{\prime\prime})\|_{2}$ $\leq C\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|}+\|M_{22}\|_{2}K_{n}r_{n}$
$\leq C\sqrt{n}\log n\times K_{n}r_{n}.$
Note that $M_{21}$ is the matrix $L_{[n]\backslash F\times(I_{s}\cup(I_{m}\backslash I_{m}^{\prime}))}$ and $M_{22}$ is the matrix $L_{[n]\backslash F\times I_{m}^{\prime}}$. We let $J_{*}$ be one of the $J_{i}$ that is in $I_{m}^{\prime}$ but has no intersection with $F$, where we will perform small ball probability and tensorization.
[($F$)$I_{s}$][($J_{*}$)$I_{m}^{\prime}\subset I_{m}$]
Figure 4. $|I_{s}|>|I_{m}|$
[($I_{s}$][$F$)$I_{m}^{\prime}\subset I_{m}$($J_{*}$)]
Figure 5. $|I_{s}|<|I_{m}|$
For each $i\in([n]\setminus F)\setminus J_{*}$, we let $X_{i}$ be the $i$th row of $\overline{L}_{n}-\lambda_{0}$. Then we can write $|X_{i}\cdot{\mathbf{v}}^{\prime}|=O(\log n\times K_{n}r_{n})$ as
$|\sum_{j\in J_{*}}x_{ij}(v_{j}^{\prime}-v_{i}^{\prime})+\sum_{j\in[n]\setminus(J_{*}\cup\{i\})}x_{ij}(v_{j}^{\prime}-v_{i}^{\prime})+f|=O(\log n\times K_{n}r_{n})$ (33)
where $v_{i}^{\prime}$ are the entries of ${\mathbf{v}}^{\prime}$, and $f$ depends on the $v_{i}^{\prime}$’s but not on the $x_{ij}$’s.
The $x_{ij}$ for $i\in([n]\setminus F)\setminus J_{*}$ and $j\in J_{*}$ are all independent. We further condition on $x_{ij}$ for $i\in([n]\setminus F)\setminus J_{*}$ and $j\not\in J_{*}$. Since $J_{*}\cap F=\emptyset$, none of the $x_{ij}$ for $i\in([N]\setminus F)\setminus J_{*}$ and $j\in J_{*}$ have been conditioned on. Thus after our conditioning, the probability of (33) holding is at most $\rho_{L,\log n\times K_{n}r_{n}}({\mathbf{v}}_{J_{*}}^{\prime}).$ We have $n-|F|-|J_{*}|$ independent such inequality. Thus the probability that, for a given ${\mathbf{v}}_{s}$ and ${\mathbf{v}}_{I_{m}\setminus{{I^{\prime}_{m}}}},$ we have the event in the claim, is at most (by Lemma 3.3)
$\rho_{L,\log n\times K_{n}r_{n}}({\mathbf{v}}_{J_{*}}^{\prime})^{n-|F|-|J_{*}|}.$
Since $J_{*}$ is one of the $J_{i}$ that is a subset of $I_{m}$, by definition of $I_{m}$ and the highest multiplicity of coordinates of ${\mathbf{v}}_{J_{*}}$ is $<|J_{*}|-\alpha_{n}\lambda_{n}n$. More precisely, by the definition of non-sparse interval (25), for any $a$ there are at least $\alpha_{n}\lambda_{n}n$ entries $i\in J_{*}$ such that (noting that over the mix part, $v_{i}=v_{i}^{\prime}$)
$|v_{i}^{\prime}-a\mathbf{1}/\sqrt{n}|\geq t_{0}^{1/2}r_{n}.$
Theorem 5.1 (more specifically (21), and also in the spirit of Remark 5.2) imply that, with $r=t_{0}^{1/2}r_{n}$ and $R=\sqrt{\log n}K_{n}r_{n}$
$\rho_{L,\log n\times K_{n}r_{n}}({\mathbf{v}}_{J_{*}}^{\prime})=O(\frac{\log n\times K_{n}}{t_{0}^{1/2}}\frac{1}{\sqrt{\alpha_{n}\lambda_{n}n}})=O(\frac{\log^{3/2}n\times\alpha_{n}^{-1/2}\lambda_{n}^{-3/2}c_{0}^{-1}(c^{\ast}_{n})^{-1}}{t_{0}^{1/2}})=O(\frac{1}{t_{0}^{1/4}}).$
Note that with the choices of parameters $\lambda_{n}$ from Theorem 2.3, $c^{\ast}_{n}$ from (31), $\alpha_{n}$ from (24) and $t_{0}$ from (23) we see the that above is much smaller than 1.
We have $|J_{*}|\leq\lambda_{n}n/2$, $|F|=|M|$. By tensorization (Lemma 3.3), the probability that, for a given ${\mathbf{v}}_{s}^{\prime}$ and ${\mathbf{v}}_{I_{m}\setminus{{I^{\prime}_{m}}}}^{\prime}$, we have the event in the claim is at most
$\left(\frac{1}{t_{0}^{1/4}}\right)^{n-|F|-\lambda_{n}n/2}\leq\left(\frac{1}{t_{0}^{1/4}}\right)^{|I_{s}|-\lambda_{n}n/2}\leq\left(\frac{1}{t_{0}^{1/4}}\right)^{\lambda_{n}n/2}.$
The total number of ${\mathbf{v}}_{s}^{\prime}$ and ${\mathbf{v}}_{I_{m}\setminus{{I^{\prime}_{m}}}}^{\prime}$ is bounded by $D^{i_{0}\alpha_{n}\lambda_{n}n}\times D\sqrt{n}$ and $D^{c^{\ast}_{n}n}$, where the first bound comes from the choices of non-sparse elements times the cardinality of net from Remark 5.3, while the second bound comes from $|I_{m}|-|I_{m}^{\prime}|=c^{\ast}_{n}n$. Note that
$D^{i_{0}\alpha_{n}\lambda_{n}n}\times D\sqrt{n},D^{c^{\ast}_{n}n}\leq n^{O(\frac{n}{\log^{2}n})}=e^{O(\frac{n}{\log n})}.$
Hence we obtain the following upper bound for the event ${\mathcal{E}}_{2.1}$ (given the choice of $t_{0}$ from (23))
$D^{i_{0}\alpha_{n}\lambda_{n}n}\times D\sqrt{n}\times D^{c^{\ast}_{n}n}\times\left(\frac{1}{t_{0}^{1/4}}\right)^{\lambda_{n}n/2}\leq\left(\frac{1}{t_{0}}\right)^{\lambda_{n}n/16}.$
The event ${\mathcal{E}}_{2}$ is the union of ${\mathcal{E}}_{drop}$ with $\mathcal{E}_{2.1}(S,{{I^{\prime}_{m}}},{{{{I^{\prime}_{s}}}}})$ over $S,{{I^{\prime}_{m}}},{{{{I^{\prime}_{s}}}}}$ with $|S|-|{{I^{\prime}_{m}}}|=c_{\ast}|I_{m}|$ and $n-|S|\geq\lambda_{n}n$ and $|S|$ a possible value of $I_{m}$. There are at most $2^{\lambda_{n}^{-1}}$ possible values of $I_{m}$ (and hence $S$). Taking union bound over all the choices, and over all $I$ (of Theorem 2.3) of size $\lambda_{n}n$ we obtain an entropy bound say at most $4^{n}$, with room to spare. We obtain the following bound for the event (29) within Case 2
$4^{n}\left(\frac{1}{t_{0}}\right)^{\lambda_{n}n/16}\leq e^{-\Theta(n)},$
given the choice of $t_{0}$ from (23) and $\lambda_{n}$ from Theorem 2.3.
6. No-structure delocalization: proof of Theorem 2.4 for the approximated vectors
We let ${\mathcal{E}}_{non-gap}$ denote the event in Theorem 2.3, where we have showed in the previous section that
${\mathbb{P}}({\mathcal{E}}_{non-gap})=1-n^{{-\omega(1)}}.$
We also let ${\mathcal{E}}_{iid-norm}$ be the event that the matrix $\overline{L}_{n}$ with zeros on the diagonal (instead of the entries of ${\operatorname{diag}}(D_{n}-{\mathbf{E}}D_{n})$) has norm $O(\sqrt{n})$. It is well known that ${\mathcal{E}}_{iid-norm}$ holds with probability $1-e^{{\Theta(n)}}$.
Throughout this section, if not specified otherwise, we will condition on these two overwhelming events ${\mathcal{E}}_{non-gap}$ and ${\mathcal{E}}_{iid-norm}$.
Theorem 2.3 applied to $\lambda_{n}=1/\log\log n$ yields that
Corollary 6.1.
Any subvectors ${\mathbf{v}}_{J}$ of ${\mathbf{v}}$ of size $|J|=n_{0}/\sqrt{\log\log n}$ satisfies
$\|{\mathbf{v}}_{J}-a1_{J}\|_{2}\geq\frac{1}{(\log n)^{A_{0}\log\log n}}=:c_{n}(=n^{{-o(1)}}).$
Assume that there exists an approximated vector ${\mathbf{v}}$ and an index $I$ such that
$\mathbf{LCD}_{\kappa,\gamma}({\mathbf{v}}_{I})\in[D_{0},2D_{0}),$ (34)
where $D_{0}\asymp n^{A}$, for which under ${\mathcal{E}}_{non-gap}$, by Theorem 2.3 and Lemma 3.5 we have that $A\geq 1/2$, for appropriate choices of $\kappa,\gamma$ (such as $\kappa=n^{1/3}$ and $\gamma=c_{n}^{3}$ – which satisfy the conditions of Lemma 3.5 with $c_{0}$ a sufficiently small constant, and with $c_{1}=1/(\log n)^{A_{0}/c_{0}}$ and say $C=10$ respectively).
The key idea to show that the event from (34) has small probability is as follows:
  • Step 1: as $\mathbf{LCD}({\mathbf{v}}_{I})$ is small, ${\mathbf{v}}_{I}$ has structure. However, as we don’t have any information on ${\mathbf{v}}_{I^{c}}$, we will use the guide discussion in Subsection 2.5 to solve for ${\mathbf{v}}_{I^{c}}$ by conditioning on some parts of the matrix $\overline{L}_{n}$. We will do so by relying on Theorem 2.6, see (41).
  • Step 2: starting from (41), which roughly can be expressed as an event $\|M_{11}{\mathbf{v}}_{I}-{\mathbf{f}}\|_{2}$ being small for some fixed ${\mathbf{f}}$, we will work with the randomness of $M_{11}$, a principle minor of size $n_{0}\times n_{0}$. As this matrix is symmetric, we have to decompose into smaller rectangular blocks to see independent random rows and entries, see Figure 7. One major problem here is that the entries on the diagonal linearly depend on other entries of the rectangular blocks, and hence we have to use the $\rho_{L}$ notion from Theorem 5.1 rather than its $\rho$ counterpart. Because there are many shifts to deal with, we will start from the shift giving worst small ball probability (see (35)). This approach is significantly different from the regularized $\mathbf{LCD}$ structures used in [Ver14] to deal with random symmetric matrices.
Toward the second step, we will need to work with some more special form of ${\mathbf{v}}_{I}$ rather than knowing that its $\mathbf{LCD}$ is large. We first fix a subset $I$ of size $n_{0}$ in $[n]$. Divide $I$ into $k=\lfloor\sqrt{\log\log n}\rfloor$ intervals $I_{1},\dots,I_{k}$ (say of indices arranged consecutively in $I$), each of size approximately
$m=n_{0}/k.$
The vector ${\mathbf{v}}_{I}$ is decomposed into ${\mathbf{v}}_{I_{1}},\dots,{\mathbf{v}}_{I_{k}}$ accordingly.
Given $A$, with sufficiently large $C$ (which is also used in the approximations of Subsection 2.4), and let
${\mathcal{A}}:=\{a\in{\mathbb{R}},|a|\leq n^{-1/2}\log^{2}n,a=\frac{l}{n^{C+1}},l\in{\mathbb{Z}}\}.$
and let
$\gamma^{\prime}=\gamma/c_{n}=c_{n}^{2},\mbox{ which has order $n^{{-o(1)}}$.}$
Define  [^6] We note that $D({\mathbf{v}}_{I})$ does depend on the partition sets $I_{1},\dots,I_{k}$ and on $\kappa,\gamma^{\prime}$, but we bypass this dependency in the notation for simplicity.
$D({\mathbf{v}}_{I}):=\max_{j}\min_{a_{j}\in{\mathcal{A}}}\mathbf{LCD}_{\kappa,\gamma^{\prime}}(\frac{{\mathbf{v}}_{I_{j}}-a\mathbf{1}}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}\|_{2}})=\max_{j}D_{j}.$ (35)
We first notice that, as $a_{j}\in{\mathcal{A}}$ and $c_{n}\leq\|{\mathbf{v}}_{I_{j}}\|_{2}\leq 1$, by Corollary 6.1
$c_{n}\leq\|{\mathbf{v}}_{I_{j}}-a_{j}\mathbf{1}\|_{2}\leq\log^{4}n.$ (36)
Claim 6.2.
Assume that $\min_{a\in{\mathcal{A}}}\mathbf{LCD}_{\kappa,\gamma^{\prime}}(\frac{{\mathbf{v}}_{I_{j}}-a\mathbf{1}}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}\|_{2}})=D_{j}$, where $D_{j}=n^{{O(1)}}$. Then for any $a\in{\mathcal{A}}$, for any $r\geq 1/D_{j}$,
${\mathbb{P}}(|X\cdot({\mathbf{v}}_{I_{j}}-a\mathbf{1})|\leq r)=O(r\log^{{O(1)}}n).$
Proof.
By definition, for any $a\in{\mathcal{A}}$ we have $\mathbf{LCD}_{\kappa,\gamma^{\prime}}(\frac{{\mathbf{v}}_{I_{j}}-a\mathbf{1}}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}\|_{2}})\geq D_{j}$. Hence by Theorem 3.7 (noting that under ${\mathcal{E}}_{non-gap}$ (36) holds), for any $r\geq 1/D_{j}\geq 1/\mathbf{LCD}_{\kappa,\gamma^{\prime}}(\frac{{\mathbf{v}}_{I_{j}}-a\mathbf{1}}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}\|_{2}})$,
$\sup_{b}{\mathbb{P}}(|X\cdot({\mathbf{v}}_{I_{j}}-a\mathbf{1})-b|\leq r)$ $=\sup_{b}{\mathbb{P}}(|X\cdot\frac{{\mathbf{v}}_{I_{j}}-a\mathbf{1}}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}}\|_{2}-b|\leq\frac{r}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}\|_{2}})$
$=O(r(\gamma^{\prime}c_{n})^{-1}+\exp(-\kappa^{2}))=O(r\log^{O(1)}n).$
Claim 6.3.
We have
$c_{n}^{-1}D({\mathbf{v}}_{I})\leq\mathbf{LCD}_{\kappa,\gamma}({\mathbf{v}}_{I}).$
Proof.
Assume $D({\mathbf{v}}_{I})=D$ is achieved at ${\mathbf{v}}_{I_{1}}$ with the corresponding shift $a_{1}\in{\mathcal{A}}$. Then by Fact 3.6, as $1\leq\|{\mathbf{v}}_{I}\|_{2}/\|{\mathbf{v}}_{I_{1}}\|_{2}\leq c_{n}^{-1}$,
$\mathbf{LCD}_{\kappa,\gamma}({\mathbf{v}}_{I})\geq c_{n}^{-1}\mathbf{LCD}_{\kappa,\gamma^{\prime}}({\mathbf{v}}_{1})\geq c_{n}^{-1}\mathbf{LCD}_{\kappa,\gamma^{\prime}}({\mathbf{v}}_{1}-a_{1}).$
As a consequence of the above claim, as $c_{n}=n^{{-o(1)}}$, instead of (34) we will be working with the event (recalling (35))
$D({\mathbf{v}}_{I})=\max_{j}D_{j}\in[D_{0},2D_{0}),$ (37)
for some $n^{1/2-o(1)}\leq D_{0}\leq n^{A}$.
Note that within this event, by definition we have every $D_{j}\ll D_{0}$, and hence we can $O(\kappa/D)$-approximate all the subvectors ${\mathbf{v}}_{j}-a_{j}\mathbf{1}$ (for some $a_{j}\in{\mathcal{A}}$). Note that the norm of each of the vector ${\mathbf{v}}_{I_{j}}-a_{j}\mathbf{1}$ is at least $c_{n}$ by (36), so by Lemma 3.8, by gluing the nets of the subvectors together, we obtain an $O(\kappa k/D_{0})$-net ${\mathcal{N}}_{I}$ of size $(C_{0}c_{n}^{-1}D_{0}/\sqrt{m})^{km}|{\mathcal{A}}|^{k}$ for the vectors ${\mathbf{v}}_{I}$, where (37) holds.
Now for a fixed ${\mathbf{v}}_{I}$ in ${\mathcal{N}}_{I}$, consider the event that there exists ${\mathbf{v}}_{I^{c}}\in{\mathbb{R}}^{n-n_{0}}$ (of norm of order 1, more specifically ${\mathbf{v}}_{I^{c}}\in\frac{1}{n^{C+1}}{\mathbb{Z}}^{n-n_{0}}$) so that
$\|(\overline{L}_{n}-\lambda_{0}){\mathbf{v}}\|_{2}\leq\sqrt{n}\log n\times\kappa k/D_{0}=:\varepsilon.$ (38)
If we let $M_{11},M_{12},M_{21}$ and $M_{22}$ be the submatrices of row and columns indexed from $I$ and $I^{c}$ (for instance $M_{22}$ is a square matrix of size $|I^{c}|=(n-n_{0})$), then we can rewrite the above as
$\|(M_{11}{\mathbf{v}}_{I},M_{21}{\mathbf{v}}_{I})+(M_{12}{\mathbf{v}}_{I^{c}},M_{22}{\mathbf{v}}_{I^{c}})\|_{2}\leq\varepsilon.$ (39)
In particularly, $\|M_{21}{\mathbf{v}}_{I}+M_{22}{\mathbf{v}}_{I^{c}}\|_{2}\leq\varepsilon$.
[…]
Figure 6. Finding non-structured subvectors.
In what follows, let
$k_{0}:=d_{n}(n-n_{0}),\mbox{ where, say }d_{n}=(\frac{1}{\log n})^{2}.$ (40)
We next apply Theorem 2.6 and Lemma 3.1 to the matrix $M_{22}$: with probability at least $1-e^{-\Theta(k_{0}^{3/2}/\log n)}$, there exists $T\subset I^{c}$ of size $|T|=n-n_{0}-k_{0}$ such that the matrix $M_{22}|_{T}$ is near isometry in the sense that
$\sigma_{n-n_{0}-k_{0}}(M_{22}|_{T})\gg d_{n}\sqrt{n-n_{0}}.$
Let $M_{22}^{\prime}$ be an $(n-n_{0}-k_{0})\times(n-n_{0})$ matrix where $M_{22}^{\prime}M_{22}|_{T}=I_{n-n_{0}-k_{0}}$ (i.e. left inverse). By the above, we can assume
$\|M_{22}^{\prime}\|_{2}=O(1/d_{n}\sqrt{n-n_{0}})=O(1/d_{n}\sqrt{n}).$
Then from $\|M_{21}{\mathbf{v}}_{I}+M_{22}{\mathbf{v}}_{I^{c}}\|_{2}\leq\varepsilon$ of (39), after being applied by $M_{22}^{\prime}$, we deduce that
$\|M_{22}^{\prime}M_{21}{\mathbf{v}}_{I}+M_{22}^{\prime}M_{22}|_{T^{c}}{\mathbf{v}}_{I^{c}}|_{T^{c}}+{\mathbf{v}}_{I^{c}}|_{T}\|_{2}\leq\varepsilon\|M_{22}^{\prime}\|_{2}\leq\varepsilon/d_{n}\sqrt{n}.$
In other words, if we fix a realization of $M_{21}$ (and hence $M_{12}$) and $M_{22}$, and if we fix $T$ and a choice of ${\mathbf{v}}_{I^{c}}|_{T^{c}}$, then ${\mathbf{v}}_{I^{c}}|_{T}$ lies within $M_{22}^{\prime}M_{21}{\mathbf{v}}_{I}+M_{22}^{\prime}M_{22}|_{T^{c}}{\mathbf{v}}_{I^{c}}|_{T^{c}}$ of distance at most $\varepsilon/d_{n}\sqrt{n}$. So if we write $M_{12}{\mathbf{v}}_{I^{c}}$ in the form $M_{12}^{\prime}{\mathbf{v}}_{I^{c}}|_{T^{c}}+M_{12}^{\prime\prime}{\mathbf{v}}_{I^{c}}|_{T}$ then we have
$\Big{\|}M_{12}{\mathbf{v}}_{I^{c}}-M_{12}^{\prime}{\mathbf{v}}_{I^{c}}|_{T^{c}}-M_{12}^{\prime\prime}(M_{22}^{\prime}M_{21}{\mathbf{v}}_{I}+M_{22}^{\prime}M_{22}|_{T^{c}}{\mathbf{v}}_{I^{c}}|_{T^{c}})\Big{\|}_{2}\leq\|M_{12}\|_{2}(\varepsilon/\sqrt{n})=O(\varepsilon/d_{n}),$
where we recall that under ${\mathcal{E}}_{iid-norm}$, $\|M_{12}\|_{2}=O(\sqrt{n})$. Together with the first half of (39), $\|M_{11}{\mathbf{v}}_{I}+M_{12}{\mathbf{v}}_{I^{c}}\|_{2}\leq\varepsilon$, by triangle inequality
$\Big{\|}M_{11}{\mathbf{v}}_{I}-M_{12}^{\prime}{\mathbf{v}}_{I^{c}}|_{T^{c}}-M_{12}^{\prime\prime}(M_{22}^{\prime}M_{21}{\mathbf{v}}_{I}+M_{22}^{\prime}M_{22}|_{T^{c}}{\mathbf{v}}_{I^{c}}|_{T^{c}})\Big{\|}_{2}=O(\varepsilon/d_{n}).$ (41)
Note that $M_{11}$ has size $n_{0}\times n_{0}$, and recall the definition of $\varepsilon$ from (38). We will bound the probability with respect to the off-diagonal entries of $M_{11}$, while other entries of $\overline{L}_{n}$ are held fix. More specifically, we can write the above as
$\|(M_{11}-F){\mathbf{v}}_{I}\|_{2}=O((\sqrt{n}\log n)\kappa k/D_{0}d_{n})=O(\sqrt{n}\rho),$
where $F$ is a deterministic matrix, and where for short we set
$\rho:=(d_{n}^{-1}\log n\times\kappa k)/D_{0}.$
[…]
Figure 7. Decomposition into i.i.d. parts.
Assume that $D(v_{I})$ is achieved at the interval $I_{j_{0}}$. Then by the rectangular decomposition trick (see Figure 7), we can pass to bounding the event
$\|(M_{11})_{(I\backslash I_{j_{0}})\times I_{j_{0}}}{\mathbf{v}}_{I_{j_{0}}}-{\mathbf{w}}\|_{2}=O(\sqrt{n}\rho)$
for some deterministic vector ${\mathbf{w}}$, where the randomness now is on the entries of $(M_{11})_{(I\backslash I_{j_{0}})}$, which are all independent.
Next, for each $i\in I\backslash I_{j_{0}}$ where $|v_{i}|\leq n^{{-1/2}}\log^{2}n$ we consider the event $|{\mathbf{r}}_{i}(M_{11})_{(I\backslash I_{j_{0}})\times I_{j_{0}}}\cdot{\mathbf{v}}_{I_{j_{0}}}-w_{i}|=O(\rho)$, which has the form
$|x_{1}(v_{i_{1}}-v_{i})+\dots+x_{m}(v_{i_{m}}-v_{i})-f|=O(\rho)$
where $x_{1},\dots,x_{m}$ are independent Bernoulli of parameter $p$, and where we recall that $km=n_{0}$ (with $k=\lfloor\sqrt{\log\log n}\rfloor$ and $n/k\leq n_{0}\leq n$), and $i_{1},\dots,i_{m}$ are the elements of $I_{j_{0}}$ and $f$ might depend on $v_{i_{j}}$ but not on the $x_{i}$.
We note that as $\sum_{i\in I_{j_{0}}}v_{i}^{2}\leq 1$, the set indices $i$ where $|v_{i}|\leq n^{{-1/2}}\log^{2}n$ is at least $n_{0}-m-n/\log^{4}n$. By Claim 6.2, as $D_{j_{0}}=D({\mathbf{v}}_{I})\in[D_{0},2D_{0}]$ and as $\rho\geq 1/D_{0}$,
${\mathbb{P}}\Big{(}|x_{1}(v_{i_{1}}-v_{i})+\dots+x_{m}(v_{i_{m}}-v_{i})-f|=O(\rho)\Big{)}=O(\rho\log^{{O(1)}}n)).$
By Lemma 3.3, we thus have
${\mathbb{P}}\Big{(}\|(M_{11})_{(I\backslash I_{j_{0}})\times I_{j_{0}}}{\mathbf{v}}_{I_{j_{0}}}-{\mathbf{w}}\|_{2}=O(\sqrt{n}\rho\Big{)}\leq(\rho\log^{{O(1)}}n)^{n_{0}-m-n/\log^{4}n}.$
Summing over $(C_{0}c_{n}^{-1}D_{0}/\sqrt{m})^{n_{0}}|{\mathcal{A}}|^{k}$ choices for the vector ${\mathbf{v}}_{I}$, over $(n^{C+1})^{d_{n}(n-n_{0})}\times\binom{n-n_{0}}{d_{n}(n-n_{0})}$ choices for the vector ${\mathbf{v}}_{I^{c}}|_{T^{c}}$, and over $\binom{n}{n_{0}}$ choices for $I$ we obtain the final union bound
$(\rho\log^{{O(1)}}n)^{n_{0}-m-n/\log^{4}n}\times(C_{0}c_{n}^{-1}D_{0}/\sqrt{m})^{n_{0}}|{\mathcal{A}}|^{k}\times(n^{C+1})^{d_{n}(n-n_{0})}\times\binom{n-n_{0}}{d_{n}(n-n_{0})}\times\binom{n-n_{0}}{d_{n}(n-n_{0})}\times\binom{n}{n_{0}}$
$\leq$ $(\kappa\log^{O(1)}n/D_{0})^{n_{0}-m-m/\log^{2C}n}\times(n^{o(1)}D_{0}/\sqrt{m})^{n_{0}}\times n^{n/\log n}\times(C^{\prime})^{n}=n^{-\Theta(n_{0})},$
provided that $\kappa=n^{c}$ with a sufficiently small constant $c$.
7. Simplicity of the spectrum: proof of Theorems 2.1 and 1.2
We first work with the centered model.
Proof.
(of Therem 2.1) Let $\overline{L}_{n}^{\prime}$ be obtained from $\overline{L}_{n}$ by changing the last two columns (and rows respectively) ${\mathbf{c}}_{n-1},{\mathbf{c}}_{n})$ of $\overline{L}_{n}$ to $({\mathbf{c}}_{n-1}+{\mathbf{c}}_{n})/\sqrt{2}$ and $(-{\mathbf{c}}_{n-1}+{\mathbf{c}}_{n})/\sqrt{2}$. We observe the followings.
Lemma 7.1.
The matrix $\overline{L}_{n}^{\prime}$ has the same spectrum as of $\overline{L}_{n}$.
Proof.
Let
$U_{2}=\begin{pmatrix}1/\sqrt{2}&-1/\sqrt{2}\\ 1/\sqrt{2}&1/\sqrt{2}\end{pmatrix}.$
Note that $\begin{pmatrix}I_{n-2}&0\\ 0&U_{2}\end{pmatrix}$ is orthogonal, and so
$\overline{L}_{n}^{\prime}=\begin{pmatrix}I_{n-2}&0\\ 0&U_{2}\end{pmatrix}^{T}\overline{L}_{n}\begin{pmatrix}I_{n-2}&0\\ 0&U_{2}\end{pmatrix}$
has the same spectrum as $\overline{L}_{n}$. ∎
Therefore, it suffices to prove Theorem 2.1 for $\overline{L}^{\prime}_{n}$.
Our next lemma is that
Lemma 7.2.
The conclusion of Theorem 2.3 and 2.4 holds for the egivenvectors of all principles $\overline{L}^{\prime}_{n-1}$ of $\overline{L}^{\prime}_{n}$ of size $(n-1)\times(n-1)$.
Proof.
The proof of this result is identical to the proofs of Theorems 2.3 and 2.4; therefore we omit the details. ∎
Let ${\mathbf{u}}=\left(\begin{array}[]{c}{\mathbf{w}}\\ b\end{array}\right)$ be an eigenvector of $\overline{L}^{\prime}_{n}$, where ${\mathbf{w}}\in{\mathbb{R}}^{n-1}$, with eigenvalue $\lambda_{i}(\overline{L}^{\prime}_{n})$. By Theorem 2.3 and Claim 3.4 we might assume that $|b|\geq n^{{-1/2-o(1)}}$ (as these results imply that with probability $1-n^{-\omega(1)}$ the set of indices where $|b|\geq n^{{-1/2-o(1)}}$ is of size $\Theta(n)$).
We consider the decomposition
$\overline{L}^{\prime}_{n}=\left(\begin{array}[]{cc}\overline{L}^{\prime}_{n-1}&{\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})\\ {\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})^{T}&f\end{array}\right).$
where ${\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})$ is a $(n-1)\times 1$ vector  [^7] We slightly abused our standard notation, here ${\mathbf{c}}_{n}$ is not exactly the last column of $\overline{L}^{\prime}_{n}$ but only a subvector. and $f\in{\mathbb{R}}$. Argue as in the discussion following Theorem 2.1, let ${\mathbf{v}}$ be a unit vector of $\overline{L}^{\prime}_{n-1}$ corresponding to $\lambda_{i}(\overline{L}^{\prime}_{n-1})$, then we have
$|{\mathbf{v}}\cdot{\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})|\leq n^{{o(1)}}\delta.$
Let $I_{n-2}$ be the set of vertices $v_{1},\dots,v_{n-2}$ which has exactly one neighbor with $v_{n}$ and $v_{n-1}$. Then the neighbor switching process changes the coordinates over $I_{n-2}$ of ${\mathbf{c}}_{n}$ randomly and independently. Furthermore, with probability $1-\exp(-\Theta(n))$, we know that
$n/4-n/8\leq|I_{n-2}|\leq n/4+n/8.$
Conditioned on the event of probability $1-n^{-\omega(1)}$ from Theorem 2.4 (applied to $\overline{L}^{\prime}_{n-1}$) that $\mathbf{LCD}_{\kappa,\gamma}({\mathbf{v}}_{I_{n-2}})\geq n^{2A}$, by Theorem 3.7, for any $\delta\geq n^{-A}$ we have that
${\mathbb{P}}_{{\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})}(|{\mathbf{v}}\cdot{\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})|\leq n^{o(1)}\delta)\leq\sup_{b}{\mathbb{P}}_{{\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})}(|{\mathbf{v}}_{I_{n-2}}\cdot{\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})_{I_{n-2}}-b|\leq n^{o(1)}\delta)=O(n^{o(1)}\delta).$
We next conclude with a proof of the non-centered model, where we recall that in our notation, the real eigenvalues of a symmetrix matrix $M_{n}$ are arranged as $\lambda_{1}(M_{n})\leq\dots\leq\lambda_{n}(M_{n})$.
Proof.
(of Therem 1.2) Now that we have established the result for $\overline{L}_{n}$, we translate the result back to $L_{n}$. Assume that the spectral decomposition of $\overline{L}_{n}$ is (with $\lambda_{1}\leq\dots\leq\lambda_{n}$)
$\overline{L}_{n}=\sum_{i}\lambda_{i}{\mathbf{v}}_{i}^{T}{\mathbf{v}}_{i},$
where one of the eigenvalue is $\lambda_{i_{0}}=0$, with ${\mathbf{v}}_{i_{0}}=\overline{\mathbf{1}}=\mathbf{1}/\sqrt{n}$, and where the eigenvectors ${\mathbf{v}}_{i}$ forms an orthonormal basis. Note that we have given effective gaps between the $\lambda_{i}$.
Observe that ${\mathbf{E}}L_{n}=-pJ_{n}+pnI_{n}$ where $J_{n}$ is the $n\times n$ matrix of ones. So
$L_{n}=\overline{L}_{n}+{\mathbf{E}}L_{n}=\sum_{i}\lambda_{i}{\mathbf{v}}_{i}^{T}{\mathbf{v}}_{i}-pJ_{n}+pnI_{n}=\sum_{i\neq i_{0}}(\lambda_{i}+pn){\mathbf{v}}_{i}^{T}{\mathbf{v}}_{i}+(pn-pn)\overline{\mathbf{1}}^{T}\overline{\mathbf{1}}.$
Since $L_{n}$ is positive semidefinite, the zero eigenvalue of $L_{n}$ is the smallest eigenvalue $\lambda_{1}(L_{n})$, with eigenvector $\overline{\mathbf{1}}$  [^8] Also, $L_{n}$ and $\overline{L}_{n}$ have the same set of eigenvectors.. So for Therem 1.2, it suffices to bound the gap of the two smallest eigenvalues of $L_{n}$, $\lambda_{2}(L_{n})-\lambda_{1}(L_{n})=\lambda_{2}(L_{n})$.
Note that $L_{n}=D_{n}-A_{n}$. We can view $A_{n}$ as a perturbation of a ${\mathbf{E}}A_{n}=pJ-pI$, which has eigenvalues $-p$ with multiplicity $n-1$ and $(p-1)n$ with multiplicity one. By classical bounds on the norm of a Wigner matrix, we have that $\|A_{n}-{\mathbf{E}}A_{n}\|=O_{p}(\sqrt{n})$ with probability $1-\exp(-\Theta(n))$. Therefore, by Weyl’s inequalities,
$\lambda_{n}(A_{n})-\lambda_{n-1}(A_{n})\geq pn/2$
with probability at least $1-\exp(-\Theta(n))$. Thus, on this event
$\lambda_{2}({\mathbf{E}}D_{n}-A_{n})-\lambda_{1}({\mathbf{E}}D_{n}-A_{n})\geq pn/2.$
Finally, by Weyl’s bound, the spectral gap $\lambda_{2}(L_{n})-\lambda_{1}(L_{n})$ of $L_{n}$ is then at least
$pn/2-2\|D_{n}-{\mathbf{E}}D_{n}\|.$
Chernoff’s bound and a simple union bound tells us that $\|D_{n}-{\mathbf{E}}D_{n}\|_{2}\leq pn/10$ with probability at least $1-\exp(-\Theta(n))$. Therefore, with probability at least $1-\exp(-\Theta(n))$, $\lambda_{2}(L_{n})\geq pn/3$, which is well above the gap size and probability proved for $L_{n}$.
8. Proof of Theorem 2.7
Proof.
(of Therem 2.7) As $\overline{L}_{n}$ and $L_{n}$ have the same set of eigenvectors, it suffices to work with $\overline{L}_{n}$. We consider the decomposition
$\overline{L}_{n}=\left(\begin{array}[]{ccc}\overline{L}_{n-2}&{\mathbf{c}}_{n-1}(\overline{L}_{n})&{\mathbf{c}}_{n}(\overline{L}_{n})\\ {\mathbf{c}}_{n-1}(\overline{L}_{n})^{T}&f&g\\ {\mathbf{c}}_{n}(\overline{L}_{n})^{T}&g&h\end{array}\right).$
where ${\mathbf{c}}_{n-1}(\overline{L}_{n})$ and ${\mathbf{c}}_{n}(\overline{L}_{n})$ are $(n-2)\times 1$ vectors  [^9] As in the previous section, here ${\mathbf{c}}_{n-1},{\mathbf{c}}_{n}$ are not exactly the last two columns of $\overline{L}_{n}$. and $f,g,h\in{\mathbb{R}}$. We slightly modify the argument following Theorem 2.1. We consider the eigenvalue equation,
$\left(\begin{array}[]{ccc}\overline{L}_{n-2}&{\mathbf{c}}_{n-1}(\overline{L}_{n})&{\mathbf{c}}_{n}(\overline{L}_{n})\\ {\mathbf{c}}_{n-1}(\overline{L}_{n})^{T}&f&g\\ {\mathbf{c}}_{n}(\overline{L}_{n})^{T}&g&h\end{array}\right)\begin{pmatrix}{\mathbf{w}}\\ a\\ b\end{pmatrix}=\lambda_{i}(\overline{L}_{n})\begin{pmatrix}{\mathbf{w}}\\ a\\ b\end{pmatrix}$ (42)
with $|a|,|b|\leq n^{-B}$ and ${\mathbf{w}}\in{\mathbb{R}}^{n-2}$. If we extract the top $n-2$ coordinates, this implies that
$\overline{L}_{n-2}{\mathbf{w}}+a{\mathbf{c}}_{n-1}(\overline{L}_{n})+b{\mathbf{c}}_{n}(L_{n})=\lambda_{i}(\overline{L}_{n}){\mathbf{w}}.$
For $B$ large enough, since we can control the size of ${\mathbf{c}}_{n-1}(\overline{L}_{n})$ and ${\mathbf{c}}_{n}(\overline{L}_{n})$, this implies that
$\|(\overline{L}_{n-2}-\lambda_{i}(\overline{L}_{n})){\mathbf{w}}\|_{2}\leq n^{-B/2}.$
In other words, ${\mathbf{w}}$ is an approximated vector of $\overline{L}_{n-2}$.
Additionally, the final coordinate of (42) yields
$|{\mathbf{c}}_{n}(\overline{L}_{n})^{T}{\mathbf{w}}|\leq|ga|+|hb|+|\lambda_{i}(\overline{L}_{n})||b|$
which for large enough $B$ implies that
$|{\mathbf{c}}_{n}(\overline{L}_{n})^{T}{\mathbf{w}}|\leq n^{-B/2}.$
The rest of the proof is almost identical to the proof of Therem 2.1. Let $I_{n-2}$ be the set of vertices $v_{1},\dots,v_{n-2}$ which has exactly one neighbor with $v_{n}$ and $v_{n-1}$. Then the neighbor switching process changes the coordinates over $I_{n-2}$ of ${\mathbf{c}}_{n}(\overline{L}_{n})$ randomly and independently. Furthermore, with probability $1-\exp(-\Theta(n))$, we know that
$n/4-n/8\leq|I_{n-2}|\leq n/4+n/8.$
Conditioned on the event of probability $1-n^{-\omega(1)}$ from Theorem 2.5 (applied to $\overline{L}_{n-2}$ and with sufficiently large $B$) that $\mathbf{LCD}_{\kappa,\gamma}({\mathbf{w}}_{I_{n-2}})\geq n^{2A}$, by Theorem 3.7, for $\delta=n^{-A-3}$ we have that
${\mathbb{P}}_{{\mathbf{c}}_{n}(\overline{L}_{n})}(|{\mathbf{w}}\cdot{\mathbf{c}}_{n}(\overline{L}_{n-2})|\leq n^{-B/2})$ $\leq{\mathbb{P}}_{{\mathbf{c}}_{n}(\overline{L}_{n})}(|{\mathbf{w}}\cdot{\mathbf{c}}_{n}(\overline{L}_{n})|\leq n^{o(1)}\delta)$
$\leq\sup_{b}{\mathbb{P}}_{{\mathbf{c}}_{n}(\overline{L}_{n})}(|{\mathbf{w}}_{I_{n-2}}\cdot{\mathbf{c}}_{n}(\overline{L_{n}})_{I_{n-2}}-b|\leq n^{o(1)}\delta)$
$=O(n^{o(1)}\delta).$
We can then take a union bound over the $\binom{n}{2}$ possible sets of pairs of coordinates.